### v1

# NIMO: a Nonlinear Interpretable MOdel

Abstract
Neural networks (NNs) have achieved tremendous success over the past decade, yet they are still extremely difficult to interpret. In contrast, linear models are less expressive but offer inherent interpretability. Linear coefficients are interpretable as the marginal effect of a feature on the prediction, assuming all other features are kept fixed. To combine the benefits of both approaches, we introduce NIMO (Nonlinear Interpretable MOdel). The key idea is to define a model where the NN is designed to learn non-linear corrections to the linear model predictions, while also maintaining the original interpretability of the linear coefficients. Relevantly, we develop an optimization algorithm based on profile likelihood that elegantly allows for optimizing over the NN parameters while updating the linear coefficients analytically. By relying on adaptive ridge regression we can easily incorporate sparsity constraints as well. We show empirically that we can recover the underlying linear coefficients while significantly improving the predictive accuracy. Compared to other hybrid interpretable approaches, our model is the only one that actually maintains the same interpretability of linear coefficients as in linear models. We also achieve higher performance on various regression and classification settings.
1 Introduction
Over the course of the past decade machine learning research has witnessed enormous success thanks to neural networks. Applications range from computer vision, natural language processing, speech recognition, robotics, finance and many more. Along with the development of these remarkable capabilities came the desire for understanding the underlying decision-making process. However, neural networks still lack a clear interpretability, which prevents their widespread use in high-stake areas like healthcare. Most research has focused on how to address this problem, either by developing intrinsically interpretable models or by providing post-hoc explanations of model predictions.
In this work, we focus on the design of an intrinsically interpretable model by employing the inherent transparency of linear models. Because of the linear relationship between the input data and the output predictions, linear models are simple and their decision-making process is transparent. The sign of a linear coefficient indicates the direction of the relationship (positive or negative), and its magnitude reflects the significance of that feature on the prediction. However, the simplicity of linear models constrains their expressiveness. In contrast, neural networks have shown great ability in learning complex feature maps from the input data in order to predict the output. While this lead to an enormous success in many applications, NNs lack interpretability of the predictions in terms of the input features, hence the term “black-box” models. This is particularly evident in the medical domain where, in order to make informed decisions, practitioners need to understand which factors lead to a particular model outcome.
To solve this dilemma, we introduce the Nonlinear Interpretable MOdel (NIMO), which inherits the expressivity of neural networks and also maintains the transparency and interpretability of linear models. More specifically, we start from a linear model and from the associated linear coefficients $\bm{\beta}$ that provide a clear interpretability at a population level. Then, we adjust each $\beta_{j}$ by a multiplicative factor $h_{j}$ according to the individual point $\bm{x}$. We learn such correction terms with a neural network $\text{NN}_{\bm{u}}$ so that the model can learn nonlinear interaction terms beyond the linear ones:
$y=\sum_{j=1}^{p}x_{j}\beta_{j}h_{j}\quad\text{with}\quad h_{j}=\text{NN}_{\bm{u}}(\bm{x})\;.$ (1)
If the neural network has a suitable design, the resulting model maintains the interpretability of linear coefficients $\bm{\beta}$ at a global level while also providing a nonlinear correction at the local level, depending on the data point $\bm{x}$. In Figure 1 we illustrate how the proposed approach shares the same high-level structure of a linear model while also integrating nonlinear corrections through neural networks. In theory we would need one neural network per feature, but in practice we can use a single network with the feature appended with positional encoding. In Section 3 we show that, with simple requirements on the neural network, we can preserve the original interpretability of the linear coefficients $\bm{\beta}$. Since the network acts as a nonlinear correction to the linear coefficients $\bm{\beta}$, the joint optimization of $\bm{\beta}$ and $\bm{u}$ might not be trivial in practice. To address this issue, we develop an optimization method based on the profile likelihood which allows for sparsity constraints as well. Specifically, we alternate a gradient step over the neural network parameters with an analytical update over the linear coefficients $\bm{\beta}$. Lastly, we show that we can extend the present framework to generalized linear models.
As commonly done in the literature, we further enforce interpretability by means of sparsity. When fewer components (features or connections) are active, it is easier to understand how the information leading to some prediction is processed. In the linear part of our model, we employ L1 regularization to enforce the sparsity in the coefficients $\bm{\beta}$. In the nonlinear part originating from the network, we resort to the group L2 regularization on the weight matrix of the first fully connected (fc) layer. We also employ a novel noise injection technique to the output of the first fc layer in order to encourage the selection of the most relevant features.
Overall, the contributions of the present work can be summarized as follows:
  • We introduce a nonlinear interpretable model (NIMO) that combines the expressive power of neural networks with the well-established interpretability of linear models.
  • We introduce an optimization algorithm that allows to jointly optimize the linear coefficients and the nonlinear contributions of the neural network.
  • On synthetic data, we show that we retrieve the underlying linear coefficients, learn the correct nonlinearities and outperform state-of-the-art hybrid models. On real data, we demonstrate NIMO’s interpretability by showing that we recover sparse coefficients while achieving performance on par with neural network approaches.
Figure 1: Comparison of linear models, neural networks, and our method. Linear models offer straightforward interpretability but limited expressivity while neural networks provide high expressivity but lack interpretability. Our method combines the strengths of both by building on the linear model and adjusting the global linear coefficients with individual nonlinear corrections.
2 Related Work
Interpretability in machine learning. Interpretability refers to the capacity to express what a model has learned and the factors influencing its outputs in a manner that is clear and understandable to humans. In the literature, the two terms interpretability and explainability are sometimes used interchangeably. However, they have subtle yet important differences. Explainability focuses on providing reasons for the model’s output. Most of the methods, especially in deep learning, provide post-hoc explanations. The most popular ones include LIME (Ribeiro et al., 2016), SHAP (Lundberg and Lee, 2017) and Grad-CAM (Selvaraju et al., 2020). On the other hand, interpretability focuses on understanding the inner mechanisms and decision processes of the model, and most of the methods involve inherently understandable models, such as decision trees (Wu et al., 2017, 2021), Lasso (Tibshirani, 1996) and Explainable Boosting Machines (Lou et al., 2013). Often, simple interpretability comes at the cost of reduced expressivity. The literature on interpretable machine learning is vast and we refer readers to (Rudin et al., 2022; Molnar, 2025) for a more in-depth review. In this work, we develop a model that has a straightforward interpretability and that, at the same time, is very expressive and flexible.
Interpretability of linear models. Linear models are inherently interpretable. They model the relationship between features and the target variable as a linear combination, making the effect of each feature easy to understand and quantify (Molnar, 2025). For example, consider a linear regression model used to predict the risk of heart disease based on a set of clinical features. Then, the coefficient associated with a specific clinical feature represents the change in the predicted risk when that feature increases by one unit, assuming all other feature values remain fixed. Methods like Lasso regularization (Tibshirani, 1996) can be further applied to achieve sparse feature selection and focus on the most relevant features. However, linear models have limited expressiveness and fail to capture complex relationships among the features. In this work, we maintain the interpretability of linear models and, at the same time, we increase the model expressiveness through a neural network.
Hybrid models. Several works have proposed to combine linear models with neural networks to achieve more interpretable models. However, none of them actually retain the original interpretability of linear coefficients as in linear models. LassoNet (Lemhadri et al., 2021) is the first approach to integrate a neural network with Lasso regression. It allows a feature to be used in the neural network only if its corresponding Lasso coefficient is nonzero. This constraint ensures that both the neural network’s nonlinear component and the Lasso linear component operate on the same subset of features. Even though this allows for feature selection, deactivating the features that are deemed unimportant in the linear part significantly constrains the capacity of the neural network. Contextual Lasso (Thompson et al., 2023) is a type of varying coefficient model (Hastie and Tibshirani, 1993). The linear regression coefficients are context-dependent and generated by a neural network. This structure provides flexibility but comes at the cost of sacrificing global interpretability of the coefficients, which is key for the interpretability of linear models. In this work, we develop an inherently interpretable model called NIMO, which maintains the global interpretability of linear models while also learning nonlinear corrections to the linear predictions.
3 Proposed approach
In Section 3.1 we introduce our non-linear interpretable model, namely NIMO, and argue why we can still retain the full interpretability of linear models while also allowing for non-linearities. Then, in Section 3.2 we propose an algorithm to elegantly train the model on the neural network parameters while analytically updating the linear coefficients. In Section 3.3 we show how to find relevant nonlinear interactions by means of sparsity. Lastly, In Section 3.4 we discuss the limitations.
3.1 A Nonlinear Interpretable Model
Model definition
Let $X\in\mathbb{R}^{n\times d}$ be $n$ $d$-dimensional observations and let $\bm{y}\in\mathbb{R}^{n}$ be the targets. In the following, we assume the data $X$ to be standardized, i.e. that each feature has zero mean and unit standard deviation. A linear model is simply defined by the linear coefficients $\bm{\beta}\in\mathbb{R}^{d}$ and a bias term $\beta_{0}\in\mathbb{R}$ as $y_{i}\approx\beta_{0}+\sum_{j=1}^{d}x_{j}\beta_{j}$. In linear models the coefficient $\beta_{j}$ is a proxy for the importance of the $j$-th feature. More precisely, the linear coefficient $\beta_{j}$ can be interpreted as the additive effect of the $j$-th feature on the output with all other variables at baseline level. The idea behind our approach is to maintain the interpretability of global linear coefficients while allowing for individual nonlinear corrections to the linear predictions. In order to accomplish this, we define our approach starting from a linear model and by multiplying the linear coefficients $\beta_{j}$ by a nonlinear correction term $h_{j}$ that depends on the individual data point $\bm{x}$. The model is defined as follows:
$f(\bm{x})=\beta_{0}+\sum_{j=1}^{d}x_{j}\beta_{j}\underbrace{(1+g_{\bm{u}_{j}}(\bm{\bm{x}_{-j}}))}_{h_{j}}\quad\text{s.t.}\quad g_{\bm{u}_{j}}(\mathbf{0})=0\;,$ (2)
where $\bm{\bm{x}_{-j}}$ is the vector $\bm{x}$ without the $j$-th component and where $g_{\bm{u}_{j}}:\mathbb{R}^{d-1}\mapsto\mathbb{R}$ are $d$ different neural networks, each parametrized by $\bm{u}_{j}$. We provide an illustration of the proposed approach in Figure 2. To make sure that our model in Eq. (2) fulfills the intended interpretability we took care of three aspects. First, the network $g_{\bm{u}_{j}}$ needs to act as a non-linear correction to the linear prediction. This is easily achieved by designing the neural network as an additive contribution to the linear prediction. We multiply the coefficient $\beta_{j}$ by $h_{j}=1+g_{\bm{u}_{j}}$, which renders the additive effect: $f(\bm{x})=\beta_{0}+\sum_{j=1}^{d}x_{j}\beta_{j}+\sum_{j=1}^{d}x_{j}\beta_{j}g_{\bm{u}_{j}}(\bm{\bm{x}_{-j}})$. Second, the neural network $g_{\bm{u}_{j}}$ cannot depend on the $j$-th feature itself. Otherwise, $x_{j}$ would contribute to the prediction not only through the linear term $\beta_{j}$ but also through the network itself, which would change the behavior and interpretation of the coefficient $\beta_{j}$. Third, we make sure that the non-linear contribution coming from the neural network keeps the feature at baseline level (as in the linear part), i.e. $g_{\bm{u}_{j}}(\bm{\bm{x}_{-j}}=\bm{0})=\bm{0}$. It is easy to see that when $\bm{\bm{x}_{-j}}=\bm{0}$ the model reduces to classical linear regression.
Figure 2: NIMO architecture. The model consists of global linear coefficients $\beta_{j}$ modulated by the individual data point $\bm{x}$ through $h_{j}$. In particular, $h_{j}$ depends on $\bm{\bm{x}_{-j}}$, i.e. all features except the $j$-th.
Model interpretability
At a high level, the proposed approach resembles a linear model where each coefficient is given by the product $\beta_{j}h_{j}$. Intuitively, $\beta_{j}$ is independent of the data point $\bm{x}$ and acts at a population level. In contrast, $h_{j}$ provides a nonlinear correction to $\beta_{j}$ and depends on the individual data point $\bm{x}$ (more precisely $\bm{\bm{x}_{-j}}$) through $g_{\bm{u}}$. This construction allows to keep the global interpretability as in a linear model while also allowing for a local nonlinear correction. To see why this is relevant, consider a healthcare application where we want to predict a treatment outcome based on some clinical data. The resulting linear coefficients represents at a population level the contribution of each feature to the prediction. However, the global explanation might not be a good fit for all patients, which is a limitation of linear models. For some patients the outcome might be better explained by a different coefficient, which we can learn by multiplying the global $\beta_{j}$ by a local $h_{j}$ specific to that patient. Overall, the proposed approach provides a population-wise interpretation of linear coefficients while still allowing for interpretable nonlinear and local corrections.
Implementation details
The proposed model in Eq. (2) requires $d$ different neural networks, which can require significant memory in high-dimensional settings. Instead, we use a single shared network $g_{\bm{u}}$ and employ positional encoding to make the network aware of which component is being queried. This can be easily done by appending the positional encoding of the $j$-th component to the input $\bm{\bm{x}_{-j}}$. In practice, $g_{\bm{u}}$ takes the full vector $\bm{x}$ as input and we can simply mask out the $j$-th component of $\bm{x}$ to get $g_{\bm{u}}(\bm{\bm{x}_{-j}})$. Furthermore, we enforce $g_{\bm{u}}(\bm{\bm{x}_{-j}}=\bm{0})=0$ by design by simply subtracting the mean prediction at each forward pass: $g_{\bm{u}}(\bm{\bm{x}_{-j}})\coloneqq g_{\bm{u}}(\bm{\bm{x}_{-j}})-g_{\bm{u}}(\bm{0})$.
3.2 Training through profile likelihood with adaptive ridge regression
For illustrative purposes, in this section we provide a high-level description of the optimization algorithm used to train NIMO. We initially consider a regression setting and later extend the same reasoning to generalized linear models, in particular for the logistic regression case.
Optimization for sparse regression
Consider the data $X\in\mathbb{R}^{n\times d}$ and targets $\bm{y}\in\mathbb{R}^{d}$. As argued before, we can model the $d$ networks with one network $g_{\bm{u}}$ and represent its outputs in a matrix $G_{\bm{u}}=g_{\bm{u}}(X)\in\mathbb{R}^{n\times d}$. The model in Eq. (2) can then be re-written in matrix form as $f(X)=B_{\bm{u}}\bm{\beta}$, where $B_{\bm{u}}=X+X\circ G_{\bm{u}}$ and $\circ$” denotes element-wise multiplication. Let us first consider the standard ridge regression setting: $\mathcal{L}(\bm{\beta},\bm{u})=\lVert\bm{y}-B_{\bm{u}}\bm{\beta}\rVert^{2}+\lambda\lVert\bm{\beta}\rVert^{2}$. To disentangle the optimization of $\bm{\beta}$ and $\bm{u}$, we take inspiration from the profile likelihood (Venzon and Moolgavkar, 1988) and consider an optimization based on $\bm{u}$ only. Since we know the analytical solution of $\bm{\beta}$ for ridge regression, we can alternate one optimization step over $\bm{u}$ with an analytical update over $\bm{\beta}$:
$\min_{\bm{\beta},\bm{u}}\mathcal{L}(\bm{\beta},\bm{u})\quad\rightarrow\quad\min_{\bm{u}}\mathcal{L}(\hat{\bm{\beta}},\bm{u})\quad\text{where}\quad\hat{\bm{\beta}}=(B_{\bm{u}}^{T}B_{\bm{u}}+\lambda I)^{-1}B_{\bm{u}}^{T}\mathbf{y}.$ (3)
In practice, $\hat{\bm{\beta}}$ depends on the current estimate of $\bm{u}$ via $B_{\bm{u}}$ so we alternate one gradient descent step over $\bm{u}$ with one analytical update over $\bm{\beta}$ according to $\hat{\bm{\beta}}$ in Eq. (3). Note that in practice this is much more effective than freely learning via gradient descent $\bm{\beta}$ alongside with $\bm{u}$. Furthermore, this formalism allows to include sparsity by simply replacing the L2 penalty with the L1 penalty:
$\min_{\bm{\beta},\bm{u}}\;\lVert\mathbf{y}-B_{\bm{u}}\bm{\beta}\rVert^{2}+\lambda\lVert\bm{\beta}\rVert_{1}\;.$ (4)
In contrast to the ridge objective, there is no closed-form solution for Lasso regression. Instead, we can use adaptive ridge regression, which is equivalent to Lasso (Grandvalet, 1998); see the Appendix for a detailed proof. By rewriting our objective as in adaptive ridge regression, we can enforce sparsity and find an analytical update for $\bm{\beta}$. The exact algorithm is shown in Alg. 1 in the Appendix.
Extension to generalized linear models
The same idea can be extended to any generalized linear model. For simplicity, consider logistic regression with L1 penalty. As before, we first get a closed-form expression for $\bm{\beta}$ and then substitute it back to optimize over the neural network parameters $\bm{u}$. What is different is that in each optimization step we use the iteratively reweighted least squares (IRLS) surrogate, which is a weighted least-squares approximation of the original problem:
$\min_{\bm{\beta},\bm{u}}\;\tfrac{1}{2}\bigl{\|}W^{1/2}(\mathbf{z}-B_{\bm{u}}\bm{\beta})\bigr{\|}^{2}+\lambda\|\bm{\beta}\|_{1}\;,$ (5)
where $W_{ii}=\sigma(\eta_{i})\bigl{[}1-\sigma(\eta_{i})\bigr{]}$ and $\bm{\eta}=B_{\bm{u}}\bm{\beta}$. The so-called working response $\bm{z}$ is computed as $\bm{z}=\bm{\eta}+W^{-1}\bigl{(}\mathbf{y}-\sigma(\bm{\eta})\bigr{)}$. The problem in Eq. (5) is also a lasso problem and we can still employ adaptive ridge regression to find a closed form solution for $\bm{\beta}$. Then we substitute $\bm{\beta}$ back and optimize over the neural networks parameters $\bm{u}$. A detailed derivation can be found in the Appendix.
3.3 Sparsity of nonlinear interactions
In the previous paragraphs we showed that linear coefficients $\bm{\beta}$ in NIMO are interpretable as in a linear model. Even though the neural network contribution is not interpretable per se, we can still identify relevant nonlinear interactions by means of sparsity. The idea is to force the first layer of the neural network to be sparse, hence to use only relevant features. To do so, we apply a group penalty on the weight matrix $W\in\mathbb{R}^{p\times d}$ of the first fully-connected (fc) layer, where $d$ is the number of input features and $p$ the hidden dimension of the first fc layer. In order to obtain sparsity at a input feature level, we need to consider neurons connected to such feature and penalize them as a group:
$\ell_{\text{group}}=\lambda_{\text{group}}\sum_{j=1}^{d}\|\mathbf{w}_{j}\|_{2}\;.$ (6)
where $\mathbf{w}_{j}=W[:,j]\in\mathbb{R}^{p}$ is the weight vector acting on the $j$-th feature. With this penalty on the first fc layer, we encourage it to select only certain input features. However, since an MLP computes a composition of functions, the subsequent fc layers can still compensate for the sparsity induced by the regularization in the first layer. Specifically, even if the first layer produces sparse outputs, subsequent layers can reweigh and transform these outputs to restore their influence on the final prediction. To mitigate this compensation effect, we inject noise into the output of the first layer. By doing so, we disrupt the deterministic signal path, making it harder for the subsequent layers to simply exploit the resulting weak, regularized, and noisy signal. As a result, the network is encouraged to focus on strong and robust signals associated with informative input features.This approach works very well in practice and allows to easily achieve sparsity. We provide further details in the Appendix.
3.4 Limitations
The proposed approach is designed to maintain the interpretability of linear models and, at the same time, to learn nonlinear corrections to linear predictions. Therefore, the underlying assumption is that the data can be described by adding nonlinearities over a linear model. Highly nonlinear datasets might not be suitable for the proposed approach. However, in such cases the question itself about interpretability would be ill-posed, at least according to our notion of interpretability. There is however one case that cannot be modeled by Eq. (2). Assume the underlying data generating process contains a self-interacting term of the kind $x_{j}f(x_{j})$, where $f$ is a linear or nonlinear function (e.g. $x\sin(x)$ or $x^{2}$). In order to make the linear coefficients interpretable we assumed $g_{\bm{u}_{j}}(\bm{\bm{x}_{-j}})$ not to depend on $x_{j}$, which clearly prevents us from learning such types of nonlinearities. However, if such feature is known to play a role in the dataset, simple feature engineering can solve the issue.
4 Experiments
In this section, we experimentally showcase the proposed approach. Firstly, in Section 4.1 we use a toy example to illustrate how the model works. Secondly, in Section 4.2 we show that across several synthetic datasets the model learns the actual underlying coefficients, the nonlinearities and the correct sparsity in the networks. We also achieve higher performance compared to other hybrid approaches. Lastly, in Section 4.3 we test the model on real datasets and show that it is able to select a sparser set of features than Lasso, while maintaining consistency with Lasso in terms of the signs and magnitudes of the selected coefficients. However, thanks to the nonlinear terms the model achieves significantly higher performance, on par with neural networks but with the benefits of inherent interpretability. In the Appendix we provide further details on each experimental setup.
Figure 3: Illustrative toy example. Left: NIMO learns the GT coefficients and also finds the correct sparsity. Middle: the learned nonlinearities coincide with the GT. Right: the first fc layer of the NN learns the correct sparsity: only features $x_{1}$ and $x_{2}$ have non-zero weights, while $x_{3}$ is ignored.
4.1 Toy example
We first illustrate NIMO on a simple three-dimensional toy example. We generate the data as follows:
$y=3\cdot x_{1}\cdot(1+\tanh(10x_{2}))+(-3)\cdot x_{2}\cdot(1+\sin(-2x_{1}))+0\cdot x_{3}+\epsilon\;,$ (7)
where $\epsilon$ is the usual Gaussian noise. This toy example has 3 features $\{x_{`},x_{2},x_{3}\}$, but $x_{3}$ is actually inactive (hence uninformative) since it is multiplied by a zero coefficient. We generate 400 samples and use a 200-100-100 train-validation-test split. As a baseline, we compare with Lasso regression. As shown in Figure 3 (left) the $\bm{\beta}$ coefficients learned with NIMO coincide with the ground truth. Relevantly, NIMO correctly learns $\beta_{3}$ to be zero, i.e. that $x_{3}$ is uninformative. In Figure 3 (middle) we show that the neural network used in NIMO can perfectly recover the underlying nonlinear interactions both for $x_{1}$ and $x_{2}$. Furthermore, if we analyze the weights of the first fc layer, we can see that the neurons connected to the input $x_{3}$ are sparse; see Figure 3 (right). This provides further interpretability in the model because it implies that $x_{3}$ does not contribute to nonlinear interactions.
4.2 Synthetic Experiments
We now showcase that NIMO is able to recover sparse coefficients and learn the underlying nonlinearities across a set of different regression and classification settings, with different number of features and sparsity levels. We compare out method against Lasso regression (Tibshirani, 1996), a vanilla neural network and two hybrid approaches, namely LassoNet (Lemhadri et al., 2021), Contextual Lasso (Thompson et al., 2023). Contextual Lasso requires the features to be separated in explanatory and contextual, which is not directly applicable to our settings since we do not have contextual features. To still compare with Contextual Lasso, we use all features as context. Compared with other hybrid approaches, NIMO achieves better or equivalent performance while still retaining interpretability. In all settings we use 200 training samples and 100 for validation and test. Further details can be found in the Appendix.
Figure 4: Learned $\bm{\beta}$ coefficients on synthetic regression datasets and comparison with Lasso regression and Ground truth. Vanilla regression (left) is purely linear while Settings 1, 2, 3 (right) have different nonlinearities. Both NIMO and Lasso recover the correct nonzero coefficients (informative features) but only NIMO recovers the correct sparsity of the zero coefficients (uninformative features).
Regression We design synthetic regression datasets by explicitly controlling the linear coefficients and the nonlinearities, similarly to Eq. (7). We explore multiple nonlinearities, different dimensionality and also use several uninformative features (i.e. with zero coefficient) to test if we can recover sparse signals. As a sanity check, we first test NIMO on a Vanilla regression dataset with linear coefficients in $\{-5,-4,...,3,4\}$ but without nonlinearities. The idea is to check if the neural network of NIMO interferes with the linear part in the absence of nonlinearities. As shown in Figure 4 (left), NIMO is able to perfectly recover the ground truth coefficients, and the neural network does not interfere with them. We provide three more regression settings (Setting 1, 2, 3) where we test different nonlinearities and different sparsity levels. We provide details about the exact data generating process in the Appendix. In Table 1 we report the MSE on the test set over all settings for NIMO and we compare it against linear, nonlinear and hybrid approaches. Overall, NIMO achieves the best performance across all settings with nonlinearities. This is possible because NIMO learns the correct coefficients (Figure 4) and, at the same time, accurately learns the underlying nonlinear interactions. Differently from Lasso, we can also perfectly learn the sparsity pattern in the coefficients, i.e. which features are uninformative; see Figure 4 (bottom, right). In contrast, nonlinear models and hybrid models learn spurious nonlinear relationships.
Table 1: MSE loss for synthetic regression settings (see the Appendix for more details). We compare with Lasso regression, NN, LassoNet and Contextual Lasso. In each setting we use 200 samples.
Features Lasso NN NIMO LassoNet Contextual Lasso
Vanilla regr. 10 0.010 0.939 0.048 0.015 -
Toy Example 3 18.982 0.446 0.166 0.628 0.646
Setting 1 5 3.164 1.109 0.030 0.078 0.381
Setting 2 10 3.340 1.482 0.217 2.991 3.418
Setting 3 50 13.122 13.718 0.334 1.342 13.070
Classification Our method can be extended to generalized linear models and here we showcase it for logistic regression. In order to create data for synthetic classification we follow a similar procedure as in the regression case, where we directly control the linear coefficients and the nonlinear interactions. The only difference is that we further apply a link function and binarize the output and get the labels. Note that this way we loose information about the original linear coefficients, which cannot be recovered exactly anymore. The sparsity in the coefficients still remains crucial. Also in this case we show various settings for different dimensionality, nonlinear terms and different number of uninformative features. We provide more details about each setting in the Appendix. In Table 2 we report the classification accuracy on the test set for all methods. Despite our efforts, we were not able to train Contextual lasso on such classification settings, even when all features were given as contextual features. Overall, our model significantly outperforms logistic regression on all settings. When analyzing the learned coefficient, which for brevity we report in the Appendix, we can see that only NIMO recovers the correct sparsity patterns. As argued before, this allows to learn the correct nonlinear interactions. Other nonlinear models and hybrid approaches perform similarly to NIMO but fail to provide the interpretability of the coefficients as in a linear model.
Table 2: Classification accuracy for synthetic classification settings (see the Appendix for more details). We compare with logistic regression, NN and LassoNet. In each setting we use 200 samples.
Features Log. Regr. NN NIMO LassoNet
Setting 1 3 0.59 0.90 0.92 0.89
Setting 2 10 0.68 0.83 0.85 0.82
Setting 3 50 0.74 0.69 0.84 0.90
Figure 5: Diabetes datasets results. Left: MSE loss on test set. All methods achieve comparable results. Middle: Learned linear coefficients by Lasso and NIMO. NIMO selects sparser coefficients (“s1, “s2”, “s4”). Right: the first fc layer of NIMO is very sparse; few nonlinearities are relevant. Errorbars are obtained by running all models across 5 different datasets splits.
4.3 Real Datasets
We further evaluated our model on two real-world datasets: the diabetes dataset (Efron et al., 2004) and the Boston housing dataset (Belsley et al., 2005). In all the experiments, we use a 60%-20%-20% train-validation-test split. We compare our model with Lasso regression, a standard neural network and LassoNet (Lemhadri et al., 2021) in terms of MSE loss and feature sparsity.
Diabetes dataset
The diabetes dataset is a classic regression dataset consisting of 442 patients, each with 10 numeric features, and the target measures disease progression one year after baseline. As shown in Figure 5 (left), we observe that all methods achieve similar MSE losses and the neural network (also the hybird methods) shows no clear advantage on this dataset. Since the diabetes dataset is commonly used for linear regression tasks, it is not surprising that the data does not exhibit complex nonlinear interactions among features. We can verify this claim by looking at the weight sparsity of the first fc layer in NIMO, shown in Figure 5 (right). Relevantly, NIMO achieves a significantly sparser solution than Lasso regression, see Figure 5 (middle). This is particularly evident by looking at the coefficients associated with the feature “s1”, “s2” and “s4”.
Boston dataset
The Boston housing dataset contains 506 instances, each with 13 features, and the target is the median value of owner-occupied homes. As shown in Figure 6, our method performs significantly better than Lasso regression and shows compatible MSE test loss with the other nonlinear and hybrid approaches. This suggests that nonlinear interactions are crucial to explain the data. While both the neural network and LassoNet are able to capture those nonlinearities, they both lack the interpretability provided by our model. It is important to note that the Boston housing dataset has been criticized for containing controversial features  [^1] https://fairlearn.org/main/user_guide/datasets/boston_housing_data.html. The feature of interest is termed “B”. From the coefficient plot in Figure 6 (middle), we observe that the significance of the feature “B” is greatly reduced in our model compared to Lasso regression. This surprising result suggests that the reason why such feature appeared to be relevant is due to some artifacts of linear models. Thanks to the decoupling between the linear coefficients and the nonlinearities in NIMO, our results suggests that the role of the feature “B” towards the prediction might be irrelevant. Note that the feature “B” is also not relevant for the nonlinear terms, as shown in Figure 6.
Figure 6: Boston datasets results. Left: Comparison of MSE losses. NIMO performs on par with other methods and significantly outperforms Lasso, suggesting nonlinear interactions are crucial. Middle: Learned linear coefficients by Lasso and NIMO. The two models learn coefficient with a comparable sparsity. Relevantly, the feature “B” is almost irrelevant for NIMO. Right: sparsity in the first fc layer of NIMO. Errorbars are obtained by running all models over 5 datasets splits.
5 Conclusions
In this paper we propose a nonlinear interpretable model. Different from other hybrid approaches, our model retains the interpretability of linear coefficients – just as in standard linear models – while also allowing for nonlinear interaction terms. However, unlike linear models, we are not limited to global linear coefficients that operate at the population level. Our model can also learn local nonlinearities, enabling the generation of modified linear coefficients tailored to individual data points. This can be viewed as an individual correction on top of the linear prediction. In order to effectively train the proposed approach, we propose an optimization algorithm based on profile likelihood and adaptive ridge regression. Specifically, we are able to optimize the neural network parameters via gradient descent while analytically updating the linear coefficients. By means of sparsity in the neural networks we can also automatically select relevant nonlinear interactions. We showcase the proposed approach on synthetic dataset and show that we can learn the underlying coefficients and nonlinearities, outperforming nonlinear and hybrid approaches. We also showcase the model on real world datasets and confirmed the same findings. When the dataset does not contain nonlinear interaction terms we recover sparse coefficients and the network becomes less relevant. When the dataset contains nonlinearities we can learn them with the same accuracy of other nonlinear approaches, but with the significant advantage of learning the interpretable linear coefficients as well.
Appendix
[…]
Appendix A Equivalence of Adaptive Ridge regression and Lasso
The equivalence in the solution of Adaptive ridge regression and Lasso regression is a well-known fact [Tibshirani, 1996, Grandvalet, 1998, Grandvalet and Canu, 1998]. Since both our model relies on this fact both in the regression and classification settings, we report its proof below.
Proof.
Let $X\in\mathbb{R}^{n\times d}$ be $n$ $d$-dimensional observations and $\bm{y}\in\mathbb{R}^{n}$ be the targets. In a linear model the relationship between target and data is assumed to be linear:
$\bm{y}=X\bm{\beta}+\epsilon\;,$ (8)
where $\bm{\beta}\in\mathbb{R}^{d}$ is the coefficient vector and $\epsilon\sim\mathcal{N}(0,1)$ is standard Gaussian noise. In Ridge Regression a penalized version of the objective is minimized:
$\min_{\bm{\beta}}\|X\bm{\beta}-\bm{y}\|_{2}^{2}+\lambda\sum_{j=1}^{d}\beta_{j}^{2}\;,$ (9)
where $\lambda$ weighs the penalty term. In Adaptive Ridge regression, instead of a single $\lambda$, a set of feature-specific penalties $\{\nu_{i}\}_{i=1}^{d}$ are introduced. The objective then reads as
$\min_{\bm{\beta}}\|X\bm{\beta}-\bm{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}\quad\text{s. t.}\quad\sum_{i=1}^{d}\frac{1}{\nu_{i}}=\frac{d}{\lambda}\;,$ (10)
where $\nu_{j}\geq 0$ $\forall j$ and $\lambda$ is a predefined value. One can solve the above constrained optimization through the method of Lagrange multipliers:
$\mathcal{L}(\bm{\nu},\mu)=\|X\bm{\beta}-\bm{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}+\mu\sum_{i=1}^{d}\frac{1}{\nu_{i}}\;,$ (11)
where $\mu$ is the Lagrangian multiplier and $\bm{\nu}\coloneqq[\nu_{1},\ldots,\nu_{d}]^{T}$. The stationary points can be found by taking the gradients with respect to $\nu_{k}$, as follows:
$\frac{\partial\mathcal{L}(\bm{\nu},\mu)}{\partial\nu_{k}}=\beta_{k}^{2}-\mu\nu_{k}^{-2}=0\quad\Rightarrow\quad\nu_{k}\beta_{k}^{2}=\mu/\nu_{k}\quad\text{or}\quad\nu_{k}|\beta_{k}|=\mu^{1/2}$ (12)
We can now use the constraint in Eq. (10):
[…]
$\Rightarrow$ $\quad\mu^{1/2}=\frac{\lambda}{d}\sum_{k}|\beta_{k}|$
If we now substitute $\nu_{k}|\beta_{k}|=\mu^{1/2}$ in the Adaptive Ridge objective in Eq. (10), we get the Lasso objective:
$f(\bm{\beta},\bm{\nu})$ $=\|X\bm{\beta}-\bm{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}$ (14)
$=\|X\bm{\beta}-\bm{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}|\beta_{i}|\cdot|\beta_{i}|$
$=\|X\bm{\beta}-\bm{y}\|^{2}+\sum_{i=1}^{d}\mu^{1/2}|\beta_{i}|$
$=\|X\bm{\beta}-\bm{y}\|^{2}+\mu^{1/2}\sum_{i=1}^{d}|\beta_{i}|.$
which gives the relationship with the lasso turns out to be a Lasso regression, with the Lasso penalty parameter at optimum to be $\mu^{1/2}=\frac{\lambda}{d}\sum_{k}|\beta_{k}|$. ∎
Appendix B Training through profile likelihood
The proposed model in Eq. (2) has two sets of parameters: the linear coefficients $\bm{\beta}$ and the neural network parameters $\bm{u}$. Since the neural network learn the nonlinear corrections over the linear predictions, the two sets of parameters are clearly entangled and in practice optimizing them both with gradient descent is tricky. Furthermore, to increase the interpretability of the model we also encourage sparse coefficients and a sparse first fc layer of the the neural network. In this section we show how to make the optimization more effective and, at the same time, how to encourage sparsity. In particular, we propose a novel optimization method based on profile likelihood with adaptive ridge regression. The profile likelihood allows for an analytical update of the linear coefficients while adaptive rigde regression provides sparsity as in Lasso regression. We further show that we can extend such an approach to generalized linear models and we showcase it in detail for logistic regression.
B.1 Regression setting: adaptive ridge regression
Let $\mathcal{L}(\bm{\beta},\bm{u})$ be the joint likelihood in Eq. (3). First, we fix the neural network parameters and analytically solve the maximization problem to obtain $\hat{\bm{\beta}}$. Then, we substitute $\hat{\bm{\beta}}$ back into the joint likelihood, which results in the profile likelihood for $\bm{u}$: $\mathcal{L}(\hat{\bm{\beta}},\bm{u})$. We can then optimize $\bm{u}$ by minimizing the negative log-profile likelihood using gradient descent. We iterate over these steps until convergence. More specifically, we draw the inspiration from the Adaptive Ridge (AR) regression and exploit its equivalence to Lasso regression [Tibshirani, 1996, Grandvalet, 1998, Grandvalet and Canu, 1998].
We already showed that the proposed model in Eq. (2) can be re-written in matrix form as:
$f(X)=B_{\bm{u}}\bm{\beta}\quad\text{with}\quad B_{\bm{u}}=X+X\circ G_{\bm{u}}$ (15)
where $\bm{\beta}\in\mathbb{R}^{d}$ are the linear coefficients, $G_{\bm{u}}\in\mathbb{R}^{n\times d}$ is the neural network output with inputs $X\in\mathbb{R}^{n\times d}$ and $\circ$” denotes element-wise multiplication. The linear model part is then expressed as:
$\bm{y}=B_{\bm{u}}\bm{\beta}+\epsilon,$ (16)
where $\bm{y}\in\mathbb{R}^{n}$ denotes the targets. Adaptive Ridge regression minimizes the following objective:
$\min_{\bm{\beta}}\|B_{\bm{u}}\bm{\beta}-\bm{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}\quad\text{s. t.}\quad\sum_{i=1}^{d}\frac{1}{\nu_{i}}=\frac{d}{\lambda}$ (17)
where $\nu_{j}>0$ are the penalty parameters for each coefficient $\beta_{j}$, and $\lambda>0$ is a predefined parameter. To avoid divergent solutions we use the re-parameterization employed in Grandvalet [1998], Grandvalet and Canu [1998]:
$\gamma_{i}=(\nu_{i}/\lambda)^{1/2}\beta_{i},\quad c_{i}=(\lambda/\nu_{i})^{1/2}\;.$ (18)
If we substitute $\beta_{i}$ and $\nu_{i}$ into the Adaptive Ridge regression objective, we get the following optimization problem:
$\min_{\bm{\gamma}}\|\bm{y}-B_{\bm{u}}D_{\bm{c}}\bm{\gamma}\|^{2}+\lambda\|\bm{\gamma}\|^{2}\quad\text{s. t.}\quad\sum_{i=1}^{d}c_{i}^{2}=d,\quad c_{i}>0,$ (19)
where $D_{\bm{c}}$ is a diagonal matrix whose diagonal elements are given by the vector $\bm{c}\coloneqq[c_{1},\ldots,c_{d}]^{T}$. With this re-parameterized problem, we can define the profile likelihood. If we fix $\bm{c}$ and $\bm{u}$, $\hat{\bm{\gamma}}$ can be analytically obtained in closed form:
$\hat{\bm{\gamma}}=\left(D_{\bm{c}}B_{\bm{u}}^{T}B_{\bm{u}}D_{\bm{c}}+\lambda I\right)^{-1}D_{\bm{c}}B_{\bm{u}}^{T}\bm{y},$ (20)
We can then get $\hat{\bm{\beta}}$ in closed form:
$\hat{\bm{\beta}}=\bm{c}\circ\hat{\bm{\gamma}}\;,$ (21)
where $\circ$” denoted element-wise multiplication. We can now substitute $\hat{\bm{\beta}}$ back into the original problem and take the constraint into consideration. This leads us to a Lagrangian minimization problem in the scaling variable $\bm{c}$ and neural network parameters $\bm{u}$:
$\min_{\bm{c},\bm{u}}\|\bm{y}-B_{\bm{u}}\hat{\bm{\beta}}\|^{2}+\tilde{\mu}\sum_{i=1}^{d}\|\bm{c}\|^{2},$ (22)
and we can optimize it through gradient descent. The training algorithm for this re-parameterized adaptive ridge regression is summarized in Alg. 1.
Algorithm 1 Training of Re-parameterized Adaptive Ridge Regression
0:  Training data $X\in\mathbb{R}^{n\times d}$, training target $\bm{y}\in\mathbb{R}^{n}$, neural network $g_{\bm{u}}(\cdot)$, regularization parameter $\lambda$, learning rate $\eta$, maximum iterations $T$.
1:  Initialize network parameters $\bm{u}$ and scaling coefficients $\bm{c}$ (with $c_{i}>0$).
2:  for $t=1,\dots,T$ do
3:     $B_{\bm{u}}=X+X\circ G_{\bm{u}}$ // Compute modified design matrix
4:     $D_{\bm{c}}=\operatorname{diag}(\bm{c})$ // Form diagonal scaling matrix
5:     $\hat{\bm{\gamma}}=\bigl{(}D_{\bm{c}}B_{\bm{u}}^{T}B_{\bm{u}}D_{\bm{c}}+\lambda I\bigr{)}^{-1}D_{\bm{c}}B_{\bm{u}}^{T}\bm{y}$ // Solve for $\hat{\bm{\gamma}}$
6:     $\hat{\bm{\beta}}=\bm{c}\circ\hat{\bm{\gamma}}$ // Compute regression coefficients
7:     $L=\|\bm{y}-B_{\bm{u}}\hat{\bm{\beta}}\|^{2}+\tilde{\mu}\sum_{i=1}^{d}\|\bm{c}\|^{2}$ // Compute loss function
8:     $\bm{c}\leftarrow\bm{c}-\eta\nabla_{\bm{c}}L,\quad\bm{u}\leftarrow\bm{u}-\eta\nabla_{\bm{u}}L$ // Gradient descent updates
9:  end for
10:  Return $\hat{\bm{\beta}},\bm{c},\bm{u}$ // Output final parameters
B.2 Logistic regression: Iteratively reweighted least squares
The procedure followed in the previous section can be extended to generalized linear models. Here we showcase how to do it for logistic regression. Logistic regression is a classification model that describes the probability that a binary outcome $y\in\{0,1\}$ occurs given a feature vector $\bm{x}\in\mathbb{R}^{d}$:
$P(y=1|\bm{x})=\sigma(\bm{x}^{T}\bm{\beta})$ (23)
where $\sigma(z)=\frac{1}{1+e^{-z}}$ is the sigmoid function. The model can then be trained by minimizing the negative log-likelihood. Similarly to the regression case discussed in the previous section, we add an adaptive ridge regression penalty to the objective:
$\mathcal{L}(\bm{\beta})=-\sum_{i=1}^{n}[y_{i}\log\sigma(\bm{x}_{i}^{T}\bm{\beta})+(1-y_{i})\log(1-\sigma(\bm{x}_{i}^{T}\bm{\beta}))]+\sum_{j=1}^{d}\nu_{j}\beta_{j}^{2}\quad\text{ s. t. }\quad\sum_{j=1}^{d}\frac{1}{\nu_{j}}=\frac{d}{\lambda}$ (24)
Now we can use the same re-parameterization used in the previous section for regression. We get the following optimization problem:
$\begin{gathered}\min_{\bm{\gamma}}-\sum_{i=1}^{n}\left[y_{i}\log\sigma\left(\sum_{j=1}^{p}x_{ij}c_{j}\gamma_{j}\right)+(1-y_{i})\log\left(1-\sigma\left(\sum_{j=1}^{p}x_{ij}c_{j}\gamma_{j}\right)\right)\right]+\lambda\|\bm{\gamma}\|^{2}\\ \text{ s.t. }\quad\sum_{j=1}^{d}c_{j}^{2}=d,\quad c_{i}>0\end{gathered}$ (25)
With the reparametrized objective we can consider the profile likelihood by fixing $\bm{c}$ and update $\bm{\gamma}$. To update $\bm{\gamma}$ we use the iteratively reweighted least squares (IRLS) algorithm. The associated update reads as:
$\bm{\gamma}^{(t+1)}=(\tilde{X}^{T}W^{(t)}\tilde{X}+\lambda I)^{-1}\tilde{X}^{T}W^{(t)}\mathbf{z}^{(t)}$ (26)
where $\tilde{X}$ is the modified design matrix with entries $\tilde{x}_{ij}=c_{j}x_{ij}$, $W^{(t)}$ is a diagonal weight matrix with entries $w_{i}^{(t)}=p_{i}^{(t)}(1-p_{i}^{(t)}),p_{i}^{(t)}=\frac{1}{1+e^{-\tilde{\bm{x}}_{i}^{T}\bm{\gamma}^{(t)}}}$, and $\mathbf{z}^{(t)}$ is the the vector of working response with entries $z_{i}^{(t)}=\bm{x}^{T}\bm{\beta}+\frac{y_{i}-p_{i}^{(t)}}{w_{i}^{(t)}}$. Note that in our model, our design matrix $X$ contains the nonlinear term as well so we would need to replace $X$ with $B_{\bm{u}}=X+X\circ G_{\bm{u}}$. We report the detailed optimization algorithm for the logistic regression case in Alg. 2.
Algorithm 2 IRLS Training for Logistic Regression
0:  Training data $X\in\mathbb{R}^{n\times d}$, target labels $\bm{y}\in\{0,1\}^{n}$, neural network $g_{\bm{u}}(\cdot)$, regularization parameters $\lambda$, learning rate $\eta$, maximum iterations $T$.
1:  Initialize parameters $\bm{u}$, scaling coefficients $\bm{c}$ (with $c_{i}>0$), and regression coefficients $\bm{\beta}$.
2:  for $t=1,\dots,T$ do
3:     $B_{\bm{u}}=X+X\circ G_{\bm{u}}$ // Compute transformed design matrix
4:     $D_{\bm{c}}=\operatorname{diag}(\bm{c})$ // Form diagonal scaling matrix
5:     $\tilde{X}=B_{\bm{u}}D_{\bm{c}}$
6:     $\hat{\bm{y}}=B_{\bm{u}}\bm{\beta}$ // Compute logits
7:     $\bm{\pi}=\sigma(\hat{\bm{y}})$
8:     $W=\operatorname{diag}(\bm{\pi}\circ(1-\bm{\pi}))$
9:     $\mathbf{z}=\hat{\bm{y}}+(\bm{y}-\bm{\pi})/\operatorname{diag}(W)$
10:     $\bm{\gamma}=(\tilde{X}^{T}W\tilde{X}+\lambda I)^{-1}\tilde{X}^{T}W\mathbf{z}$
11:     $\bm{\beta}=\bm{c}\circ\hat{\bm{\gamma}}$ // Update $\bm{\beta}$
12:     $L=\operatorname{BCE}(\hat{\bm{y}},\bm{y})+\lambda\sum_{i}c_{i}^{2}$ // Compute total loss
13:     $\bm{c}\leftarrow\bm{c}-\eta\nabla_{\bm{c}}L,\quad\bm{u}\leftarrow\bm{u}-\eta\nabla_{\bm{u}}L$ // Gradient descent step
14:  end for
15:  Return $\bm{\beta},\bm{c},\bm{u}$
B.3 Extension to sub-$\ell_{1}$ pseudo-norms
In some cases sparsity is a crucial aspect in the learning process. One of the limitations of lasso is the well-known over-shrinkage effect, which means that shrinkage of uninformative features could be achieved at the expense of shrinking the retained features as well. One way to avoid this is to use sub-$\ell_{1}$ pseudo-norms, which in the limist of the $\ell_{0}$ pseudo-norm achieve sparsity without any shrinkage. The main challenge with sub-$\ell_{1}$ pseudo-norm is that the resulting objective function becomes non-convex and harder to optimize. However, recent work has successfully used sub-$\ell_{1}$ norms as a prior on regression problems [Negri et al., 2023], which is similar to our setting.
We now show how to use a sub-$\ell_{1}$ prior in our Adaptive Ridge regression objective, which we report below for completeness:
$\min_{\bm{\beta}}\|X\bm{\beta}-\bm{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}\quad\text{s.t.}\quad\sum_{i=1}^{d}\frac{1}{\nu_{i}}=\frac{d}{\lambda}$ (27)
In order to generalize to sub-$\ell_{1}$ pseudo-norms we need the following simple modification on $\nu_{i}$:
$\sum_{i=1}^{d}\frac{1}{\nu^{\delta}}=C,\quad 0<\delta\leq 1.$ (28)
We now showcase that we can follow a procedure completely similar to what we had with the original adaptive ridge regression. Fist, we reformulate the objective with the method of Lagrangian multipliers:
$\mathcal{L}(\bm{\nu},\mu)=\|X\bm{\beta}-\bm{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}+\mu\sum_{i=1}^{d}\frac{1}{\nu_{i}^{\delta}}$ (29)
Then, we find stationary points for $\nu_{i}$ by taking the gradients and setting them to zero:
$\frac{\partial\mathcal{L}(\bm{\nu},\mu)}{\partial\nu_{i}}=\beta_{i}^{2}-\mu\delta\nu_{i}^{-(\delta+1)}=0\quad\Rightarrow$ $\quad\nu_{i}^{\delta+1}=\frac{\mu\delta}{\beta_{i}^{2}}$ (30)
$\Rightarrow$ $\quad\nu_{i}=\left(\frac{\mu\delta}{\beta_{i}^{2}}\right)^{\frac{1}{\delta+1}}$
The penalty in the original adaptive ridge regression can be rewritten as:
$\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}=\sum_{i=1}^{d}\left(\frac{\mu\delta}{\beta_{i}^{2}}\right)^{\frac{1}{\delta+1}}\beta_{i}^{2}\propto\sum_{i=1}\beta_{i}^{2(1-\frac{1}{\delta+1})}.$ (31)
Note that when $\delta=1$, we recover the Lasso $\ell_{1}$ norm. When $\delta\rightarrow 0$, it approaches the $\ell_{0}$ pseudo-norm, which basically counts the number of nonzero entries. We now apply the same re-parameterization:
$\gamma_{i}=(\nu_{i}/\lambda)^{1/2}\beta_{i},\quad c_{i}=(\lambda/\nu_{i})^{1/2}\;.$ (32)
The objective function then reads as
$\min_{\bm{\gamma}}\|\bm{y}-XD_{\bm{c}}\bm{\gamma}\|^{2}+\lambda\|\bm{\gamma}\|^{2}\quad\text{s.t.}\quad\sum_{i=1}^{d}c_{i}^{2\delta}=\lambda^{\delta}C,\quad c_{i}>0\;.$ (33)
For each fixed $\bm{c}$, $\bm{\gamma}$ can now be obtained in closed form:
$\hat{\bm{\gamma}}=\left(D_{\bm{c}}X^{T}XD_{\bm{c}}+\lambda I\right)^{-1}D_{\bm{c}}X^{T}\bm{y},$ (34)
and $\bm{\beta}$ can be computed as before by:
$\hat{\bm{\beta}}=\bm{c}\circ\hat{\bm{\gamma}}$ (35)
Once we compute the $\hat{\bm{\beta}}$, we can substitute it back into the original objective. Taking the constraint into consideration leads us to a Lagrangian minimization problem in the scaling variable $\bm{c}$ only:
$\min_{\bm{c}}\|\bm{y}-X\hat{\bm{\beta}}\|^{2}+\tilde{\mu}\sum_{i=1}^{d}c_{i}^{2\delta},$ (36)
and we can optimize it through gradient descent.
Appendix C NIMO: Implementation details
C.1 Model Formulation
We formulate NIMO as follow:
$f(\bm{x})=\beta_{0}+\sum_{j=1}^{d}x_{j}\beta_{j}(1+g_{\bm{u}_{j}}(\bm{x}_{-j}))\quad\text{s.t.}\quad g_{\bm{u}_{j}}(\mathbf{0})=0\;.$ (37)
The architecture of our model involves two parts, as shown in Fig. 2. The left part is the nonlinear neural network, which operates on the input feature $\bm{x}$ element-wise. The right part is the linear model, with coefficients $\beta_{j}$ multiplied by the neural network output and then summed together.
In practice, each feature component $x_{j}$ of the input vector $\bm{x}$ is temporarily masked out to construct $\bm{x}_{-j}$, which is then passed through a neural network $g_{\bm{u}_{j}}$. The output of this network is added to a constant $1$ and then multiplied by the original feature value $x_{j}$, effectively producing a rescaled version of that feature:
$\tilde{x}_{j}=x_{j}\cdot(1+g_{\bm{u}_{j}}(\bm{x}_{-j}))$ (38)
This rescaling can be interpreted as an element-wise modulation of the input features based on context. The rescaled features are then linearly combined using fixed coefficients $\bm{\beta}$ to produce the final prediction:
$y=\sum_{j=1}^{d}\tilde{x}_{j}\beta_{j}.$
C.2 Practical Implementation
We implement our model with PyTorch Lightning [Falcon and The PyTorch Lightning team, 2019]. While in principle each feature $x_{j}$ could have its own dedicated network $g_{\bm{u}_{j}}$, this can be computationally expensive for high-dimensional inputs. To reduce overhead, we instead use a shared neural network $g_{\bm{u}}$ across all components.
Positional Encoding
To ensure the network remains aware of which feature is currently masked out, we append a positional encoding to $\bm{x}_{-j}$ before passing it through the shared network. We assign each masked input feature $\bm{x}_{-j}$ a unique binary vector corresponding to its index $j$, represented in fixed-length binary form using $\lfloor\log_{2}d\rfloor+1$ bits. This compact encoding allows the shared neural network to remain index-aware with minimal overhead.
Group Penalty and Noise Injection
The shared network is implemented with 3 fully-connected (fc) layers. To encourage the network to select only certain relevant features contributing to the nonlinearity, we apply a group penalty on the first fc layer. However, due to the compensation effect from the subsequent layers, the sparsity is difficult to achieve by simply applying this penalty. To mitigate this issue, we inject noise into the output of the first layer by adding scaled Gaussian noise. Concretely, we draw noise from a normal distribution with mean 0 and variance 1, scale it by 0.2, and add it to the output. The noise-injected output is then passed through a tanh activation function before being fed into the subsequent layers.
Other Details
After the second fc layer, a sin activation function is applied to capture potentially repeating or high-frequency cyclic patterns in the data [Sitzmann et al., 2020]. The learning rate is 1e-3 for synthetic experiments, and 5e-3 for real-world datasets. We choose Adam [Kingma, 2014] as the default optimizer. For regression, we regularize the network also by restricting the output range within $[-1,1]$, which works well in practice. For classification, we use a less restrictive reduction to the range $[-1,3]$.
Appendix D Experiments
We conduct experiments on both synthetic and real-world datasets. All experiments are conducted on a workstation equipped with a single user-grade NVIDIA GeForce RTX 2080 Ti GPU. The methods we compare with are linear model (Lasso for regression, Logistic Regression for classification), neural network, LassoNet [Lemhadri et al., 2021] and Contextual Lasso [Thompson et al., 2023] (for regression only). Here we provide the details about the experiment setups.
D.1 Implementation details / Hyperparameters
Linear Model
We use scikit-learn [Pedregosa et al., 2011] to implement the Lasso and Logistic regression model. More specifically, we use LassoCV and LogisticRegressionCV with default 5-fold cross validation to select the best model. For LogisticRegressionCV, the $\ell_{1}$ penalty is applied.
Neural Network
We implement the naive neural network (MLP) with PyTorch Lightning [Falcon and The PyTorch Lightning team, 2019]. For a fair comparison, the MLP has the same structure as the nonlinear part in our NIMO model, except no positional encoding. We use Adam [Kingma, 2014] as the default optimizer. The learning rate is 1e-3 for synthetic experiments, and 5e-3 for real-world datasets. Strong dropout regularization (p=0.6) is used to avoid overfitting.
LassoNet
We use the original implementation of LassoNet from their Github repository  [^2] https://github.com/lasso-net/lassonet. Using the same API as shown in the LassoNet document, we configure the model to have a structure and number of parameters comparable to our model. All other hyperparameters are kept at their default values. The best model is automatically selected along the entire regularization path, as demonstrated in their paper. The one with the smallest validation loss is chosen.
Contextual Lasso
We get the code of Contextual Lasso from the original Github repository  [^3] https://github.com/ryan-thompson/ContextualLasso.jl, which was implemented in Julia language. We follow the examples in the documentation and configure the model to have a structure and number of parameters comparable to our own model.
D.2 Synthetic experiments: Setup
We create the synthetic datasets by first sampling the features, and then compute the targets according to different nonlinear patterns. Random Gaussian noises are added to the targets. The feature values are sampled uniformly between $-2$ and $2$. For logistic regression, we pass the targets through a logistic (sigmoid) function to get probabilities between 00 and $1$, and then sample the final binary labels using a Bernoulli distribution with those probabilities. In all the settings, we use $200$ training samples and $100$ samples each for validation and test.
Settings for regression
All features are indexed starting from 1. $\epsilon\sim\mathcal{N}(0,0.1^{2})$ is a Gaussian noise.
  • Setting 0 (toy example), $n=200,p=3$
    $y=$ $+3\cdot x_{1}\cdot[1+\tanh(10x_{2})]$ (39)
    $-3\cdot x_{2}\cdot\left[1+\sin(-2x_{1})\right]$
    $+\epsilon$
  • Setting 1, $n=200,p=5$
    $y=$ $+3\cdot x_{1}\cdot\left[1+(2\sigma(x_{2}x_{3})-1)\right]$ (40)
    $-2\cdot x_{2}$
    $+2\cdot x_{3}$
    $+\epsilon$
  • Setting 2, $n=200,p=10$
    $y=$ $+1\cdot x_{1}\cdot[1+\tanh(x_{2}x_{3}+\sin(x_{4}))]$ (41)
    $+2\cdot x_{2}\cdot[1+\sin(2x_{1})]$
    $-1\cdot x_{3}\cdot\left[1+\frac{2}{\pi}\arctan(x_{2}x_{4})\right]$
    $+\epsilon$
  • Setting 3, $n=200,p=50$
    $y=$ $-2\cdot x_{1}\cdot[1+\tanh(x_{2}x_{4})]$ (42)
    $+2\cdot x_{2}\cdot\left[1+\frac{2}{\pi}\arctan(x_{4}-x_{5})\right]$
    $+3\cdot x_{4}\cdot\left[1+\tanh(x_{2}+\sin(x_{5}))\right]$
    $-1\cdot x_{5}\cdot\left[1+(2\sigma(x_{1}x_{4})-1)\right]$
    $+\epsilon$
  • Setting 4, $n=200,p=10$
    $y$ $=-5\cdot x_{1}-4\cdot x_{2}-3\cdot x_{3}-\dots+0\cdot x_{6}$ (43)
    $+1\cdot x_{7}+2\cdot x_{8}+\dots+4\cdot x_{10}$
    $+\epsilon$
Settings for classification
For all the settings, we generate $y$ first, and then generate labels through the following procedure:
$\pi=\frac{1}{1+e^{-y}},\quad\text{label}\sim\text{Bernoulli}(\pi)$
  • Setting 1, $n=200,p=3$
    $y=$ $+2\cdot x_{1}\cdot[1+2\tanh(x_{2})]$ (44)
    $-2\cdot x_{2}\cdot[1+3\sin(2x_{1})+\tanh(2x_{1})]$
    $+1+\epsilon$
  • Setting 2, $n=200,p=10$
    $y=$ $+10\cdot x_{1}\cdot[1+2\cdot\tanh(2x_{2})+\sin(x_{4})]$ (45)
    $+20\cdot x_{2}\cdot[1+2\cdot\cos(2x_{1})]$
    $-20\cdot x_{3}\cdot[1+2\cdot\arctan(x_{2}x_{4})]$
    $+10\cdot x_{4}$
    $-10+\epsilon$
  • Setting 3, $n=200,p=50$
    $y=$ $-20\cdot x_{1}\cdot[1+\tanh(x_{2}x_{4})]$ (46)
    $+20\cdot x_{2}\cdot\left[1+\frac{2}{\pi}\arctan(x_{4}-x_{5})\right]$
    $+30\cdot x_{4}\cdot\left[1+\tanh(x_{2}+\sin(x_{5}))\right]$
    $-10\cdot x_{5}\cdot\left[1+(2\sigma(x_{1}x_{4})-1)\right]$
    $+\epsilon$
D.3 Synthetic experiments: Additional results
Below, we present additional experimental results for the synthetic settings described above.
Figure 7: Sparsity and feature selection in the first fc layer of NIMO for regression settings. Left: Setting 1. Middle: Setting 2. Right: Setting 3.
Regression
The advantage of our model is that it not only recovers the underlying linear coefficients but also selects sparse features for the neural network, thanks to the noise injection and group penalty. In the main paper, we presented the recovered linear coefficients. Here, we show the sparse features selected by the neural network in Figure 7. We observe that even though all features are input to the neural network, only a small subset of features is selected. Moreover, these selected features align with those involved in the nonlinearity computation in our settings, which demonstrates the model’s ability to effectively identify and focus on the most relevant features for the task.
Classification
In classification, the presence of the link function can obscure the underlying signal, so while a linear decision boundary may suffice, recovering the true generative coefficients is more difficult than in regression, where the model directly learns to approximate the target surface. Across all settings, the goal is not to recover the precise magnitudes of the ground truth coefficients, but to accurately identify their sparsity pattern and directional effects (i.e., signs). We compare the recovered coefficients from logistic regression and our model across three classification settings, as shown in Figure 8. Since recovering the exact magnitudes of the coefficients is not feasible, we normalize all coefficients, including the ground truth, by dividing them by the maximum absolute value among the coefficients, to enable a fair comparison.
Figure 8: Learned coefficients for synthetic classification settings. We compare NIMO and Logistic regression with the ground truth. Since we use binarized labels it is impossible to retrieve the exact coefficients. Instead, we compare normalized coefficients, such that the maximum is always 1.
We observe that logistic regression struggles to identify the relevant features and their correct signs. In contrast, our model generally succeeds in selecting the correct features, and the neural network component effectively captures sparse but important features that contribute to nonlinearity. It is worth noting that NIMO also selects some noisy features, especially in Setting 3. However, our experiments are primarily intended as a proof of concept to demonstrate that our model can be easily extended to generalized linear models. With careful hyperparameter tuning, NIMO is capable of selecting more accurate features. The sparse features selected by the neural network are shown in Figure 9.
Figure 9: Sparsity and feature selection in the first fc layer of NIMO for classification settings. Left: Setting 1. Middle: Setting 2. Right: Setting 3.
D.4 Real-dataset experiments
We conduct experiments on two real-world datasets: the diabetes dataset [Efron et al., 2004] and the Boston housing dataset [Belsley et al., 2005]. Each dataset is randomly split into training (60%), validation (20%), and test (20%) sets, repeated over 5 independent runs. For each split, we perform a grid search to select the optimal penalty parameter based on the validation loss (or validation accuracy for classification tasks). Error bars and confidence intervals are reported across these different splits.

### v2

@@
 Abstract
-Neural networks (NNs) have achieved tremendous success over the past decade, yet they are still extremely difficult to interpret.
-In contrast, linear models are less expressive but offer inherent interpretability.
-Linear coefficients are interpretable as the marginal effect of a feature on the prediction, assuming all other features are kept fixed.
-To combine the benefits of both approaches, we introduce NIMO (Nonlinear Interpretable MOdel).
-The key idea is to define a model where the NN is designed to learn non-linear corrections to the linear model predictions, while also maintaining the original interpretability of the linear coefficients.
-Relevantly, we develop an optimization algorithm based on profile likelihood that elegantly allows for optimizing over the NN parameters while updating the linear coefficients analytically.
-By relying on adaptive ridge regression we can easily incorporate sparsity constraints as well.
-We show empirically that we can recover the underlying linear coefficients while significantly improving the predictive accuracy.
-Compared to other hybrid interpretable approaches, our model is the only one that actually maintains the same interpretability of linear coefficients as in linear models.
-We also achieve higher performance on various regression and classification settings.
+Deep learning has achieved remarkable success across many domains, but it has also created a growing demand for interpretability in model predictions.
+Although many explainable machine learning methods have been proposed, post-hoc explanations lack guaranteed fidelity and are sensitive to hyperparameter choices, highlighting the appeal of inherently interpretable models.
+For example, linear regression provides clear feature effects through its coefficients.
+However, such models are often outperformed by more complex neural networks (NNs) that usually lack inherent interpretability.
+To address this dilemma, we introduce NIMO, a framework that combines inherent interpretability with the expressive power of neural networks.
+Building on the simple linear regression, NIMO is able to provide flexible and intelligible feature effects.
+Relevantly, we develop an optimization method based on parameter elimination, that allows for optimizing the NN parameters and linear coefficients effectively and efficiently.
+By relying on adaptive ridge regression we can easily incorporate sparsity as well.
+We show empirically that our model can provide faithful and intelligible feature effects while maintaining good predictive performance.
 1 Introduction
-Over the course of the past decade machine learning research has witnessed enormous success thanks to neural networks.
-Applications range from computer vision, natural language processing, speech recognition, robotics, finance and many more.
-Along with the development of these remarkable capabilities came the desire for understanding the underlying decision-making process.
-However, neural networks still lack a clear interpretability, which prevents their widespread use in high-stake areas like healthcare.
-Most research has focused on how to address this problem, either by developing intrinsically interpretable models or by providing post-hoc explanations of model predictions.
-In this work, we focus on the design of an intrinsically interpretable model by employing the inherent transparency of linear models.
-Because of the linear relationship between the input data and the output predictions, linear models are simple and their decision-making process is transparent.
-The sign of a linear coefficient indicates the direction of the relationship (positive or negative), and its magnitude reflects the significance of that feature on the prediction.
-However, the simplicity of linear models constrains their expressiveness.
-In contrast, neural networks have shown great ability in learning complex feature maps from the input data in order to predict the output.
-While this lead to an enormous success in many applications, NNs lack interpretability of the predictions in terms of the input features, hence the term “black-box” models.
-This is particularly evident in the medical domain where, in order to make informed decisions, practitioners need to understand which factors lead to a particular model outcome.
-To solve this dilemma, we introduce the Nonlinear Interpretable MOdel (NIMO), which inherits the expressivity of neural networks and also maintains the transparency and interpretability of linear models.
-More specifically, we start from a linear model and from the associated linear coefficients $\bm{\beta}$ that provide a clear interpretability at a population level.
-Then, we adjust each $\beta_{j}$ by a multiplicative factor $h_{j}$ according to the individual point $\bm{x}$.
-We learn such correction terms with a neural network $\text{NN}_{\bm{u}}$ so that the model can learn nonlinear interaction terms beyond the linear ones:
-$y=\sum_{j=1}^{p}x_{j}\beta_{j}h_{j}\quad\text{with}\quad h_{j}=\text{NN}_{\bm{u}}(\bm{x})\;.$ (1)
-If the neural network has a suitable design, the resulting model maintains the interpretability of linear coefficients $\bm{\beta}$ at a global level while also providing a nonlinear correction at the local level, depending on the data point $\bm{x}$.
-In Figure 1 we illustrate how the proposed approach shares the same high-level structure of a linear model while also integrating nonlinear corrections through neural networks.
-In theory we would need one neural network per feature, but in practice we can use a single network with the feature appended with positional encoding.
-In Section 3 we show that, with simple requirements on the neural network, we can preserve the original interpretability of the linear coefficients $\bm{\beta}$.
-Since the network acts as a nonlinear correction to the linear coefficients $\bm{\beta}$, the joint optimization of $\bm{\beta}$ and $\bm{u}$ might not be trivial in practice.
-To address this issue, we develop an optimization method based on the profile likelihood which allows for sparsity constraints as well.
-Specifically, we alternate a gradient step over the neural network parameters with an analytical update over the linear coefficients $\bm{\beta}$.
-Lastly, we show that we can extend the present framework to generalized linear models.
-As commonly done in the literature, we further enforce interpretability by means of sparsity.
-When fewer components (features or connections) are active, it is easier to understand how the information leading to some prediction is processed.
-In the linear part of our model, we employ L1 regularization to enforce the sparsity in the coefficients $\bm{\beta}$.
-In the nonlinear part originating from the network, we resort to the group L2 regularization on the weight matrix of the first fully connected (fc) layer.
-We also employ a novel noise injection technique to the output of the first fc layer in order to encourage the selection of the most relevant features.
-Overall, the contributions of the present work can be summarized as follows:
-  • We introduce a nonlinear interpretable model (NIMO) that combines the expressive power of neural networks with the well-established interpretability of linear models.
-  • We introduce an optimization algorithm that allows to jointly optimize the linear coefficients and the nonlinear contributions of the neural network.
-  • On synthetic data, we show that we retrieve the underlying linear coefficients, learn the correct nonlinearities and outperform state-of-the-art hybrid models.
-On real data, we demonstrate NIMO’s interpretability by showing that we recover sparse coefficients while achieving performance on par with neural network approaches.
+Over the past decade, neural networks have achieved remarkable success across domains, from computer vision (Krizhevsky et al., 2012) to natural language processing (Vaswani et al., 2017).
+At the same time, their deployment in high-stakes domains such as healthcare has created a pressing demand for interpretability (Esteva et al., 2017).
+However, neural networks in general lack interpretability of the predictions in terms of the input features, hence the term “black-box” models.
+In contrast, classical models such as linear regression and decision trees are often considered as highly interpretable because of their transparent structure and simple decision rules, but they can lack predictive power on complex, high-dimensional, or highly nonlinear tasks (Hastie, 2009; Breiman et al., 2017).
+This tension between accuracy and interpretability has motivated much recent work on models and methods that combine the expressive power of neural networks with the transparency of simpler models (Lemhadri et al., 2021; Thompson et al., 2023; Wu et al., 2017; 2021).
+Meanwhile, many model-agnostic, post-hoc explainers have been proposed, such as SHAP (Lundberg & Lee, 2017) and LIME (Ribeiro et al., 2016), providing instance-level feature attributions for arbitrary predictors.
+But these post-hoc explanations are approximations and may depend on the choices of parameters, such as the background distribution and kernel width.
+Among various ways to achieve interpretability, feature effects (Scholbeck et al., 2019) provide a direct, quantitative description of how each feature influences the prediction, something post-hoc methods can approximate but do not guarantee.
+We use the term feature effects in a broad sense, referring to how changes in input features influence model predictions.
+A common formalization of this idea is through marginal effects (Nguyen, 2020), which define feature effects as the direction and magnitude of the change in prediction due to an infinitesimal change in a feature value.
+Linear models offer immediately intelligible marginal effects through their coefficients (Molnar, 2025), but are outperformed by more complex nonlinear models, particularly neural networks.
+Unfortunately, these more expressive models typically lose the property of providing such inherently intelligible effects.
+In such models, marginal effects must instead be obtained by computing derivatives of the fitted function with respect to the inputs, usually through computationally expensive automatic differentiation (autograd) frameworks (Paszke et al., 2017).
+To bridge this gap, we propose an intrinsically interpretable hybrid model that preserves intelligible marginal effects while retaining the expressivity of neural networks.
+Specifically, we start from a linear model $y=\sum_{j}x_{j}\beta_{j}$, and adjust each $\beta_{j}$ by a multiplicative factor $h_{j}$, which is the output of a neural network $\text{NN}_{\bm{u}_{j}}(\mathbf{x})$ evaluated at the input instance $\mathbf{x}$:
+$y=\sum_{j=1}^{p}x_{j}\beta_{j}h_{j}(\mathbf{x})\quad\text{with}\quad h_{j}(\mathbf{x})=\text{NN}_{\bm{u}_{j}}(\mathbf{x})\;.$ (1)
+In Figure 1, we illustrate how the proposed approach retains the high-level structure of a linear model while integrating per-instance nonlinear corrections through neural networks.
+Within this framework, we distinguish between local explanations and global interpretations.
+Local explanations describe how changes in a feature influence the prediction for a specific instance, enabled here by the per-instance corrections $h_{j}(\mathbf{x})$ from the neural network.
+Global interpretations, in contrast, summarize the overall influence of a feature across the dataset, which we later formalize through the marginal effects at the mean (MEM).
+In our model, these coincide exactly with the global linear coefficients.
+To illustrate, consider a medical application where we aim to estimate a patient’s risk of developing a disease.
+A local explanation would ask: given this patient’s age, weight, and other characteristics, how does a small change in age affect their individual risk?
+A global interpretation instead asks: if all other variables are held fixed, how does age in general influence disease risk across the population?
+For linear models, these two perspectives coincide, but in our framework they differ due to potential nonlinear feature interactions.
+Our network is designed so that the model preserves the classical interpretation of the linear coefficients $\bm{\beta}$, while simultaneously providing per-instance corrections through $h_{j}(\mathbf{x})$.
+This enables us to unify global summaries via MEM with local explanations at the instance level within a single coherent framework.
 Figure 1: Comparison of linear models, neural networks, and our method.
@@
 Our method combines the strengths of both by building on the linear model and adjusting the global linear coefficients with individual nonlinear corrections.
+Since the network parameters $\bm{u}$ and the linear coefficients $\bm{\beta}$ are tightly coupled, jointly optimizing them is non-trivial.
+To address this, we develop an optimization method based on parameter elimination.
+Specifically, we derive a closed-form expression for the linear coefficients $\bm{\beta}$ as a function of the NN parameters $\bm{u}$.
+Substituting this expression back into the objective reduces the problem to optimizing only over the NN parameters.
+To further enhance interpretability, we impose sparsity by formulating the problem as adaptive ridge regression (Grandvalet, 1998).
+In the network, we further apply group $\ell_{2}$ regularization (Yuan & Lin, 2006) to the weight matrix of the first fully connected layer, encouraging feature-level sparsity and clarifying how inputs contribute to predictions.
+Overall, the contributions of the present work can be summarized as follows:
+  • We introduce a nonlinear interpretable model that combines the expressivity of neural networks with the interpretability of linear models, providing intelligible global interpretability via marginal effects at the mean (MEM) and per-instance corrections through the network.
+  • We introduce an optimization algorithm via parameter elimination that effectively and efficiently optimizes the linear coefficients and the neural network parameters .
+  • We systematically evaluate NIMO on synthetic and real datasets, verifying that it recovers correct feature effects, achieves accurate per-instance explanations, and offers a favorable trade-off between interpretability and predictive performance compared to existing methods.
 2 Related Work
@@
 Most of the methods, especially in deep learning, provide post-hoc explanations.
-The most popular ones include LIME (Ribeiro et al., 2016), SHAP (Lundberg and Lee, 2017) and Grad-CAM (Selvaraju et al., 2020).
-On the other hand, interpretability focuses on understanding the inner mechanisms and decision processes of the model, and most of the methods involve inherently understandable models, such as decision trees (Wu et al., 2017, 2021), Lasso (Tibshirani, 1996) and Explainable Boosting Machines (Lou et al., 2013).
-Often, simple interpretability comes at the cost of reduced expressivity.
-The literature on interpretable machine learning is vast and we refer readers to (Rudin et al., 2022; Molnar, 2025) for a more in-depth review.
-In this work, we develop a model that has a straightforward interpretability and that, at the same time, is very expressive and flexible.
-Interpretability of linear models.
-Linear models are inherently interpretable.
-They model the relationship between features and the target variable as a linear combination, making the effect of each feature easy to understand and quantify (Molnar, 2025).
-For example, consider a linear regression model used to predict the risk of heart disease based on a set of clinical features.
-Then, the coefficient associated with a specific clinical feature represents the change in the predicted risk when that feature increases by one unit, assuming all other feature values remain fixed.
-Methods like Lasso regularization (Tibshirani, 1996) can be further applied to achieve sparse feature selection and focus on the most relevant features.
-However, linear models have limited expressiveness and fail to capture complex relationships among the features.
-In this work, we maintain the interpretability of linear models and, at the same time, we increase the model expressiveness through a neural network.
+The most popular ones include LIME (Ribeiro et al., 2016), SHAP (Lundberg & Lee, 2017) and Grad-CAM (Selvaraju et al., 2020).
+On the other hand, interpretability focuses on understanding the inner mechanisms and decision processes of the model, and most of the methods that claimed to be interpretable involve inherently understandable models, such as decision trees (Wu et al., 2017; 2021), Lasso (Tibshirani, 1996) and Explainable Boosting Machines (Lou et al., 2013).
+The literature on interpretable machine learning is vast and we refer readers to Rudin et al.
+(2022) and Molnar (2025) for a more in-depth review.
+Linear Models and Input Feature Effects.
+Coefficients from linear regression models provide inherently interpretable feature effects, directly quantifying how changes in input features influence predictions (Nguyen, 2020; Molnar, 2025).
+This transparency makes linear models especially attractive in domains where interpretability is critical.
+Regularization methods such as the Lasso (Tibshirani, 1996) further enhance interpretability by selecting the most relevant input features.
+However, linear models cannot capture nonlinearities or interactions without complicated feature engineering, e.g., basis expansion, which limits their ability to reflect complex data-generating processes.
+This motivates our approach, which preserves the interpretability of linear coefficients while extending their expressiveness to capture more complex data patterns.
 Hybrid models.
-Several works have proposed to combine linear models with neural networks to achieve more interpretable models.
-However, none of them actually retain the original interpretability of linear coefficients as in linear models.
-LassoNet (Lemhadri et al., 2021) is the first approach to integrate a neural network with Lasso regression.
-It allows a feature to be used in the neural network only if its corresponding Lasso coefficient is nonzero.
-This constraint ensures that both the neural network’s nonlinear component and the Lasso linear component operate on the same subset of features.
-Even though this allows for feature selection, deactivating the features that are deemed unimportant in the linear part significantly constrains the capacity of the neural network.
-Contextual Lasso (Thompson et al., 2023) is a type of varying coefficient model (Hastie and Tibshirani, 1993).
-The linear regression coefficients are context-dependent and generated by a neural network.
-This structure provides flexibility but comes at the cost of sacrificing global interpretability of the coefficients, which is key for the interpretability of linear models.
-In this work, we develop an inherently interpretable model called NIMO, which maintains the global interpretability of linear models while also learning nonlinear corrections to the linear predictions.
+Several works have proposed combining linear models with neural networks to improve interpretability.
+Neural Additive Models (NAMs) (Agarwal et al., 2021) extend generalized additive models by learning one neural network per feature, whose outputs are summed to form the prediction.
+This design preserves interpretability at the feature level but does not have the ability to capture feature interactions.
+LassoNet (Lemhadri et al., 2021) integrates a neural network with Lasso regression by enforcing that a feature can be used in the nonlinear part only if its corresponding linear coefficient is nonzero.
+Interpretable Mesomorphic Networks (IMN) (Kadra et al., 2024) use a hypernetwork to predict linear coefficients on a per-instance basis.
+While this increases flexibility, it sacrifices global interpretability of the coefficients, making it difficult to recover clear baseline effects.
+In contrast, our proposed model (NIMO) preserves the global interpretability of linear coefficients while augmenting them with nonlinear corrections.
+This unifies global summaries via MEM with instance-level explanations within a single coherent framework.
 3 Proposed approach
-In Section 3.1 we introduce our non-linear interpretable model, namely NIMO, and argue why we can still retain the full interpretability of linear models while also allowing for non-linearities.
-Then, in Section 3.2 we propose an algorithm to elegantly train the model on the neural network parameters while analytically updating the linear coefficients.
-In Section 3.3 we show how to find relevant nonlinear interactions by means of sparsity.
-Lastly, In Section 3.4 we discuss the limitations.
 3.1 A Nonlinear Interpretable Model
 Model definition
-Let $X\in\mathbb{R}^{n\times d}$ be $n$ $d$-dimensional observations and let $\bm{y}\in\mathbb{R}^{n}$ be the targets.
+Let $X\in\mathbb{R}^{n\times d}$ be $n$ observations with $d$ dimensions and let $\mathbf{y}\in\mathbb{R}^{n}$ be the targets.
 In the following, we assume the data $X$ to be standardized, i.e.
 that each feature has zero mean and unit standard deviation.
-A linear model is simply defined by the linear coefficients $\bm{\beta}\in\mathbb{R}^{d}$ and a bias term $\beta_{0}\in\mathbb{R}$ as $y_{i}\approx\beta_{0}+\sum_{j=1}^{d}x_{j}\beta_{j}$.
-In linear models the coefficient $\beta_{j}$ is a proxy for the importance of the $j$-th feature.
-More precisely, the linear coefficient $\beta_{j}$ can be interpreted as the additive effect of the $j$-th feature on the output with all other variables at baseline level.
-The idea behind our approach is to maintain the interpretability of global linear coefficients while allowing for individual nonlinear corrections to the linear predictions.
-In order to accomplish this, we define our approach starting from a linear model and by multiplying the linear coefficients $\beta_{j}$ by a nonlinear correction term $h_{j}$ that depends on the individual data point $\bm{x}$.
-The model is defined as follows:
-$f(\bm{x})=\beta_{0}+\sum_{j=1}^{d}x_{j}\beta_{j}\underbrace{(1+g_{\bm{u}_{j}}(\bm{\bm{x}_{-j}}))}_{h_{j}}\quad\text{s.t.}\quad g_{\bm{u}_{j}}(\mathbf{0})=0\;,$ (2)
-where $\bm{\bm{x}_{-j}}$ is the vector $\bm{x}$ without the $j$-th component and where $g_{\bm{u}_{j}}:\mathbb{R}^{d-1}\mapsto\mathbb{R}$ are $d$ different neural networks, each parametrized by $\bm{u}_{j}$.
+A linear regression model is simply defined as $y=\beta_{0}+\sum_{j=1}^{d}x_{j}\beta_{j}$.
+In linear models the coefficient $\beta_{j}$ can be interpreted as the additive effect of the $j$-th feature on the output when all other feature values remain fixed.
+The idea of our model is to maintain the intelligible global interpretation at the population level while allowing for local explanations at the per-instance level.
+To do so, we define our approach starting from a linear model and multiplying the linear coefficients $\beta_{j}$ by a nonlinear correction term $h_{j}$ that depends on the data point $\mathbf{x}$:
+$f(\mathbf{x})=\beta_{0}+\sum_{j=1}^{d}x_{j}\beta_{j}\underbrace{(1+g_{\bm{u}_{j}}(\mathbf{\mathbf{x}_{-j}}))}_{h_{j}}\quad\text{s.t.}\quad g_{\bm{u}_{j}}(\mathbf{0})=0\;,$ (2)
+where $\mathbf{\mathbf{x}_{-j}}$ is the vector $\mathbf{x}$ without the $j$-th component and $g_{\bm{u}_{j}}(\cdot)$ are $d$ different scalar-valued functions defined by neural networks, each parametrized by $\bm{u}_{j}$.
 We provide an illustration of the proposed approach in Figure 2.
-To make sure that our model in Eq.
-(2) fulfills the intended interpretability we took care of three aspects.
-First, the network $g_{\bm{u}_{j}}$ needs to act as a non-linear correction to the linear prediction.
-This is easily achieved by designing the neural network as an additive contribution to the linear prediction.
-We multiply the coefficient $\beta_{j}$ by $h_{j}=1+g_{\bm{u}_{j}}$, which renders the additive effect: $f(\bm{x})=\beta_{0}+\sum_{j=1}^{d}x_{j}\beta_{j}+\sum_{j=1}^{d}x_{j}\beta_{j}g_{\bm{u}_{j}}(\bm{\bm{x}_{-j}})$.
-Second, the neural network $g_{\bm{u}_{j}}$ cannot depend on the $j$-th feature itself.
-Otherwise, $x_{j}$ would contribute to the prediction not only through the linear term $\beta_{j}$ but also through the network itself, which would change the behavior and interpretation of the coefficient $\beta_{j}$.
-Third, we make sure that the non-linear contribution coming from the neural network keeps the feature at baseline level (as in the linear part), i.e.
-$g_{\bm{u}_{j}}(\bm{\bm{x}_{-j}}=\bm{0})=\bm{0}$.
-It is easy to see that when $\bm{\bm{x}_{-j}}=\bm{0}$ the model reduces to classical linear regression.
+Due to the specific design of our model, it exhibits the following properties.
+First, by explicitly excluding the $j$-th feature as input of the neural network, i.e.
+$g_{\bm{u}_{j}}(\mathbf{\mathbf{x}_{-j}})$, $x_{j}$ contributes to the prediction only through the linear term $\beta_{j}$, preserving the behavior and interpretation of the coefficient $\beta_{j}$.
+Second, as we assume the data is standardized, the mean value of the features is zero.
+Therefore, since by construction $g_{\bm{u}_{j}}(\mathbf{0})=0$, when all other features are fixed at the mean value (i.e.
+at zero), the model reduces to a linear model.
+In other words, the marginal effect of feature $j$ is exactly $\beta_{j}$, as in a purely linear model.
+This ensures that the nonlinear correction term does not alter the global interpretability of the linear coefficients.
 Figure 2: NIMO architecture.
-The model consists of global linear coefficients $\beta_{j}$ modulated by the individual data point $\bm{x}$ through $h_{j}$.
-In particular, $h_{j}$ depends on $\bm{\bm{x}_{-j}}$, i.e.
+The model consists of global linear coefficients $\beta_{j}$ modulated by the individual data point $\mathbf{x}$ through $h_{j}$.
+In particular, $h_{j}$ depends on $\mathbf{\mathbf{x}_{-j}}$, i.e.
 all features except the $j$-th.
-Model interpretability
-At a high level, the proposed approach resembles a linear model where each coefficient is given by the product $\beta_{j}h_{j}$.
-Intuitively, $\beta_{j}$ is independent of the data point $\bm{x}$ and acts at a population level.
-In contrast, $h_{j}$ provides a nonlinear correction to $\beta_{j}$ and depends on the individual data point $\bm{x}$ (more precisely $\bm{\bm{x}_{-j}}$) through $g_{\bm{u}}$.
-This construction allows to keep the global interpretability as in a linear model while also allowing for a local nonlinear correction.
-To see why this is relevant, consider a healthcare application where we want to predict a treatment outcome based on some clinical data.
-The resulting linear coefficients represents at a population level the contribution of each feature to the prediction.
-However, the global explanation might not be a good fit for all patients, which is a limitation of linear models.
-For some patients the outcome might be better explained by a different coefficient, which we can learn by multiplying the global $\beta_{j}$ by a local $h_{j}$ specific to that patient.
-Overall, the proposed approach provides a population-wise interpretation of linear coefficients while still allowing for interpretable nonlinear and local corrections.
 Implementation details
-The proposed model in Eq.
-(2) requires $d$ different neural networks, which can require significant memory in high-dimensional settings.
-Instead, we use a single shared network $g_{\bm{u}}$ and employ positional encoding to make the network aware of which component is being queried.
-This can be easily done by appending the positional encoding of the $j$-th component to the input $\bm{\bm{x}_{-j}}$.
-In practice, $g_{\bm{u}}$ takes the full vector $\bm{x}$ as input and we can simply mask out the $j$-th component of $\bm{x}$ to get $g_{\bm{u}}(\bm{\bm{x}_{-j}})$.
-Furthermore, we enforce $g_{\bm{u}}(\bm{\bm{x}_{-j}}=\bm{0})=0$ by design by simply subtracting the mean prediction at each forward pass: $g_{\bm{u}}(\bm{\bm{x}_{-j}})\coloneqq g_{\bm{u}}(\bm{\bm{x}_{-j}})-g_{\bm{u}}(\bm{0})$.
-3.2 Training through profile likelihood with adaptive ridge regression
-For illustrative purposes, in this section we provide a high-level description of the optimization algorithm used to train NIMO.
-We initially consider a regression setting and later extend the same reasoning to generalized linear models, in particular for the logistic regression case.
+A naive implementation of Eq.
+2 requires $d$ separate neural networks, which is impractical in high-dimensional settings.
+To address this, we use a single shared network $g_{\bm{u}}$ and provide it with positional information to distinguish between features.
+Concretely, we append a positional encoding of the index $j$ to the masked input vector $\mathbf{\mathbf{x}_{-j}}$, where the $j$-th component of $\mathbf{x}$ is set to zero.
+This way, $g_{\bm{u}}$ can learn feature-specific corrections while reusing parameters across all features.
+Finally, to enforce the interpretability constraint $g_{\bm{u}}(\mathbf{0})=0$, we simply subtract the mean prediction in each forward pass: $g_{\bm{u}}(\mathbf{\mathbf{x}_{-j}})\coloneqq g_{\bm{u}}(\mathbf{\mathbf{x}_{-j}})-g_{\bm{u}}(\bm{0})$.
+3.2 Interpretability via marginal effects
+Following Nguyen (2020), the marginal effect (ME) of feature $j$ at input $\mathbf{x}$ is defined as
+$\mathrm{ME}_{j}=\frac{\partial f(\mathbf{x})}{\partial x_{j}},$ (3)
+for any model $f(\mathbf{x})$.
+By definition, $\mathrm{ME}_{j}$ measures the instantaneous change in prediction when varying $x_{j}$.
+However, marginal effects in nonlinear models are not constant, as they depend on the other input features.
+One way to summarize them is by computing marginal effects at the mean (MEM), which estimates marginal effects at the average values of the input features:
+$\mathrm{MEM}_{j}=\left.\frac{\partial f(\mathbf{x})}{\partial x_{j}}\right|_{\mathbf{\mathbf{x}}=\bar{\mathbf{x}}}.$ (4)
+Based on the definition, we can compute the marginal effects for NIMO as below:
+$\mathrm{ME}_{j}$ $=\beta_{j}\big(1+g_{\bm{u}_{j}}(\mathbf{x}_{-j})\big)+\sum_{i\neq j}x_{i}\beta_{i}\frac{\partial g_{\bm{u}_{i}}(\mathbf{x}_{-i})}{\partial x_{j}}\;,$
+$\mathrm{MEM}_{j}$ $=\beta_{j}\;.$
+Note that the mean value of input features $\bar{\mathbf{x}}=\mathbf{0}$ after standardization.
+Due to the design of our model, $\mathrm{MEM}_{j}$ coincides with the linear coefficient.
+In principle, such marginal effects can also be computed via automatic differentiation for other differentiable models, but these do not necessarily yield interpretable results.
+In contrast, for NIMO, the MEMs directly correspond to the linear coefficients $\bm{\beta}$, capturing the additive population-level contribution of each feature.
+3.3 Training via parameter elimination
+For illustrative purposes, in this section we provide a high-level description of the optimization algorithm used to train NIMO. More detailed training procedure can be found in Appendix B.
 Optimization for sparse regression
-Consider the data $X\in\mathbb{R}^{n\times d}$ and targets $\bm{y}\in\mathbb{R}^{d}$.
+Consider the data $X\in\mathbb{R}^{n\times d}$ and targets $\mathbf{y}\in\mathbb{R}^{n}$.
 As argued before, we can model the $d$ networks with one network $g_{\bm{u}}$ and represent its outputs in a matrix $G_{\bm{u}}=g_{\bm{u}}(X)\in\mathbb{R}^{n\times d}$.
 The model in Eq.
-(2) can then be re-written in matrix form as $f(X)=B_{\bm{u}}\bm{\beta}$, where $B_{\bm{u}}=X+X\circ G_{\bm{u}}$ and $\circ$” denotes element-wise multiplication.
-Let us first consider the standard ridge regression setting: $\mathcal{L}(\bm{\beta},\bm{u})=\lVert\bm{y}-B_{\bm{u}}\bm{\beta}\rVert^{2}+\lambda\lVert\bm{\beta}\rVert^{2}$.
-To disentangle the optimization of $\bm{\beta}$ and $\bm{u}$, we take inspiration from the profile likelihood (Venzon and Moolgavkar, 1988) and consider an optimization based on $\bm{u}$ only.
-Since we know the analytical solution of $\bm{\beta}$ for ridge regression, we can alternate one optimization step over $\bm{u}$ with an analytical update over $\bm{\beta}$:
-$\min_{\bm{\beta},\bm{u}}\mathcal{L}(\bm{\beta},\bm{u})\quad\rightarrow\quad\min_{\bm{u}}\mathcal{L}(\hat{\bm{\beta}},\bm{u})\quad\text{where}\quad\hat{\bm{\beta}}=(B_{\bm{u}}^{T}B_{\bm{u}}+\lambda I)^{-1}B_{\bm{u}}^{T}\mathbf{y}.$ (3)
-In practice, $\hat{\bm{\beta}}$ depends on the current estimate of $\bm{u}$ via $B_{\bm{u}}$ so we alternate one gradient descent step over $\bm{u}$ with one analytical update over $\bm{\beta}$ according to $\hat{\bm{\beta}}$ in Eq.
-(3).
-Note that in practice this is much more effective than freely learning via gradient descent $\bm{\beta}$ alongside with $\bm{u}$.
-Furthermore, this formalism allows to include sparsity by simply replacing the L2 penalty with the L1 penalty:
+2 can then be re-written in matrix form as $f(X)=B_{\bm{u}}\bm{\beta}$, where $B_{\bm{u}}=X+X\circ G_{\bm{u}}$ and $\circ$” denotes element-wise multiplication.
+Let us first consider the standard ridge regression setting: $\mathcal{L}(\bm{\beta},\bm{u})=\lVert\mathbf{y}-B_{\bm{u}}\bm{\beta}\rVert^{2}+\lambda\lVert\bm{\beta}\rVert^{2}$.
+To separate the optimization over $\bm{\beta}$ and $\bm{u}$, we take inspiration from the profile likelihood approach (Venzon & Moolgavkar, 1988; Murphy & Van der Vaart, 2000).
+The key idea is to eliminate $\bm{\beta}$ by first solving for its closed-form expression in terms of $\bm{u}$, and then substituting it back into the objective.
+This reduces the problem to an optimization over $\bm{u}$ only:
+$\min_{\bm{\beta},\bm{u}}\mathcal{L}(\bm{\beta},\bm{u})\;\;\longrightarrow\;\;\min_{\bm{u}}\mathcal{L}(\hat{\bm{\beta}}(\bm{u}),\bm{u}),\quad\text{where}\quad\hat{\bm{\beta}}(\bm{u})=(B_{\bm{u}}^{T}B_{\bm{u}}+\lambda I)^{-1}B_{\bm{u}}^{T}\mathbf{y}.$ (5)
+We can then efficiently optimize the objective with gradient descent over $\bm{u}$ only.
+Furthermore, within this framework we can incorporate sparsity by replacing the $\ell_{2}$ penalty with the $\ell_{1}$ penalty:
 $\min_{\bm{\beta},\bm{u}}\;\lVert\mathbf{y}-B_{\bm{u}}\bm{\beta}\rVert^{2}+\lambda\lVert\bm{\beta}\rVert_{1}\;.$ (6)
-In contrast to the ridge objective, there is no closed-form solution for Lasso regression.
-Instead, we can use adaptive ridge regression, which is equivalent to Lasso (Grandvalet, 1998); see the Appendix for a detailed proof.
-By rewriting our objective as in adaptive ridge regression, we can enforce sparsity and find an analytical update for $\bm{\beta}$.
-The exact algorithm is shown in Alg.
-1 in the Appendix.
+Unlike ridge regression, Lasso does not admit a closed-form solution for $\bm{\beta}$.
+To eliminate $\bm{\beta}$ in this setting, we use adaptive ridge regression, which admits at each step a closed-form expression for $\bm{\beta}$ in terms of $\bm{u}$ and is equivalent to Lasso at the optimum (Grandvalet, 1998); see Appendix A for a detailed proof.
+This enables us to efficiently optimize the problem using gradient descent while still enforcing sparsity.
+The full procedure is summarized in Algorithm 1 in the Appendix.
+A well-known limitation of Lasso is the over-shrinkage effect, where large coefficients are excessively penalized (Fan & Li, 2001).
+In some applications, this can lead to significant bias and degrade model performance.
+To mitigate this issue, we can extend the $\ell_{1}$ norm regularization to a sub-$\ell_{1}$ pseudo-norm, which penalizes large coefficients less aggressively while still promoting sparsity.
+A detailed introduction to this approach is provided in Appendix B.3.
 Extension to generalized linear models
-The same idea can be extended to any generalized linear model.
-For simplicity, consider logistic regression with L1 penalty.
-As before, we first get a closed-form expression for $\bm{\beta}$ and then substitute it back to optimize over the neural network parameters $\bm{u}$.
-What is different is that in each optimization step we use the iteratively reweighted least squares (IRLS) surrogate, which is a weighted least-squares approximation of the original problem:
+The same idea can be extended to any generalized linear model of the form: $f(\mathbb{E}[Y|\mathbf{x}])=\beta_{0}+\beta_{1}x_{1}+\dots+\beta_{p}x_{p}=\mathbf{x}^{T}\bm{\beta}$, where $f$ is a link function.
+For simplicity, consider logistic regression with $\ell_{1}$ penalty.
+As before, we first get a closed-form expression for $\bm{\beta}$ in terms of $\bm{u}$ and then substitute it back to optimize over the neural network parameters $\bm{u}$.
+What is different is that in each optimization step we use the iteratively reweighted least squares (IRLS) surrogate (Green, 1984; McCullagh & Nelder, 1989), which is a weighted least-squares approximation of the original problem:
 $\min_{\bm{\beta},\bm{u}}\;\tfrac{1}{2}\bigl\|W^{1/2}(\mathbf{z}-B_{\bm{u}}\bm{\beta})\bigr\|^{2}+\lambda\|\bm{\beta}\|_{1}\;,$ (7)
@@
 The so-called working response $\bm{z}$ is computed as $\bm{z}=\bm{\eta}+W^{-1}\bigl(\mathbf{y}-\sigma(\bm{\eta})\bigr)$.
-The problem in Eq.
-(5) is also a lasso problem and we can still employ adaptive ridge regression to find a closed form solution for $\bm{\beta}$.
-Then we substitute $\bm{\beta}$ back and optimize over the neural networks parameters $\bm{u}$.
-A detailed derivation can be found in the Appendix.
-3.3 Sparsity of nonlinear interactions
-In the previous paragraphs we showed that linear coefficients $\bm{\beta}$ in NIMO are interpretable as in a linear model.
+Eq.
+7 is also a Lasso problem and we can employ adaptive ridge regression to find a closed form expression for $\hat{\bm{\beta}}(\bm{u})$.
+Then, we substitute $\hat{\bm{\beta}}(\bm{u})$ back and optimize over the neural networks parameters $\bm{u}$ via gradient descent.
+Appendix B.2 includes a detailed derivation.
+3.4 Proof-of-concept: A toy example
+We illustrate NIMO on a simple three-dimensional toy example. The data is generated as:
+$y=3\cdot x_{1}\cdot(1+\tanh(10x_{2}))+(-3)\cdot x_{2}\cdot(1+\sin(-2x_{1}))+0\cdot x_{3}+\epsilon\;,$ (8)
+where $\epsilon\sim\mathcal{N}(0,0.1^{2})$ is Gaussian noise.
+This example involves three features $\{x_{1},x_{2},x_{3}\}$, but $x_{3}$ is inactive since it is multiplied by a zero coefficient.
+We generate 400 samples and use a 200-100-100 train-validation-test split.
+As a baseline, we compare against Lasso regression.
+We further apply group $\ell_{2}$ regularization to the weight matrix of the first fully connected layer in the neural network to enhance interpretability.
+Further implementation details are given in Appendix C.2.
+The results are shown in Figure 3.
+On the left, we see that the estimated coefficients $\bm{\beta}$, which also represent the marginal effects at the mean, learned by NIMO align with the ground truth; in particular, NIMO correctly identifies $\beta_{3}=0$, i.e.
+that $x_{3}$ is uninformative.
+In the middle, the nonlinear corrections $g_{j}(\mathbf{x}_{-j})$ recover the true underlying interactions for both $x_{1}$ and $x_{2}$.
+This highlights an important interpretability aspect: while $\beta_{1}=3$ and $\beta_{2}=-3$ represent the global population-level effects, the neural network outputs vary across individual inputs, providing per-instance corrections.
+Therefore, NIMO not only preserves the global interpretability of the coefficients but also captures the correct local deviations at the per-instance level.
+Finally, the right panel shows that neurons connected to $x_{3}$ in the first fully connected layer are sparse, permitting an additional layer of interpretability: $x_{3}$ does not contribute to nonlinear interactions.
+Figure 3: Illustrative toy example.
+Left: NIMO learns the GT coefficients and also finds the correct sparsity.
+Middle: the learned nonlinearities coincide with the GT.
+Right: the first fc layer of the NN learns the correct sparsity: only features $x_{1}$ and $x_{2}$ have non-zero weights, while $x_{3}$ is ignored.
+4 Experiments
+In this section, we experimentally evaluate the proposed NIMO, focusing on interpretability through marginal effects.
+We first introduce the baseline methods and the experimental protocol used for comparison.
+Next, we formulate three hypotheses and present empirical results that examine both the interpretability and predictive performance of NIMO.
+Additional results are provided in Appendix D.
+4.1 Methods and protocol
+Methods:
+In this paper, we compare our method against Lasso (Tibshirani, 1996), a vanilla neural network (NN(Goodfellow et al., 2016), and several state of the art interpretable approaches, namely LassoNet (Lemhadri et al., 2021), Neural Additive Models (NAMs(Agarwal et al., 2021) and Interpretable Mesomorphic Networks (IMNs(Kadra et al., 2024).
+LassoNet encourages the network to use only a subset of the available input features by enforcing sparsity with a novel objective.
+NAMs learn a linear combination of neural networks that each attend to a single input feature.
+IMNs optimize deep hypernetworks to generate explainable linear models on a per-instance basis.
+Table 1: MSE loss for synthetic regression settings.
+Method Setting 1 Setting 2 Setting 3
+(#features) (5) (10) (50)
+Lasso 3.164 3.340 13.122
+NN 1.109 1.482 13.718
+NAM 3.427 5.126 16.543
+IMN 0.137 1.188 6.308
+LassoNet 0.078 2.612 1.738
+NIMO 0.024 0.197 0.380
+Protocol:
+We conduct experiments on both synthetic and real datasets to verify the interpretability and performance of our model.
+For synthetic datasets, we create them by explicitly controlling the linear coefficients and the nonlinearities, similarly to Eq.
+8.
+We explore multiple nonlinearities, different dimensionalities and also use several uninformative features (i.e.
+with zero coefficient) to test if we can recover sparse signals.
+The details about the exact data generating process are provided in the Appendix D.2.
+For real datasets, we choose several popular and well-studied datasets from UCI Machine Learning Repository  [^1] https://archive.ics.uci.edu/datasets.
+4.2 Hypotheses and experiment results
+Hypothesis 1:
+NIMO outperforms other methods on synthetic datasets and remains robust in low-data regime.
+We compare NIMO with the above methods on the synthetic datasets and evaluate predictive performance using mean squared error (MSE).
+The results are presented in Table 1.
+In each setting, we use 200 samples for training.
+In this low-data regime, the naive neural network tends to overfit even with a relatively shallow MLP, while the linear model is too simple to capture the complex underlying nonlinearities.
+NAM, which assigns a separate MLP to each feature component, is even more prone to overfitting than the naive neural network.
+Overall, the results show that NIMO not only outperforms the other methods by effectively capturing complex nonlinear feature interactions, but also remains robust in low-data settings due to the regularization and optimization strategies employed.
+Hypothesis 2:
+NIMO accurately estimates the marginal effects at the mean (MEM), preserving the additive contribution of each feature at the population-level.
+Figure 4: Learned $\bm{\beta}$ coefficients on synthetic regression datasets and comparison with Lasso regression and Ground truth.
+Vanilla regression (left) is purely linear while Settings 1, 2, 3 (right) have different nonlinearities.
+Both NIMO and Lasso recover the correct nonzero coefficients (informative features) but only NIMO recovers the correct sparsity of the zero coefficients (uninformative features).
+This experiment evaluates whether NIMO can recover population-level interpretability in terms of marginal effects at the mean.
+For the other methods that rely on automatic differentiation, MEM can certainly be computed, but they do not yield meaningful interpretations.
+They reduce to arbitrary numerical values that neither explain the prediction for an instance nor capture population-level effects.
+In contrast, the specific structure of NIMO ensures that the marginal effects at the mean coincide with the linear coefficients, as in the linear models, providing a clear and intelligible interpretation.
+Therefore, we restrict the comparison to Lasso only, which directly yields globally interpretable coefficients.
+As a sanity check, we first test NIMO on a Vanilla regression dataset with purely linear coefficients in $\{-5,-4,\ldots,3,4\}$ and no nonlinearities.
+This verifies whether the neural network component of NIMO interferes with the linear part when no nonlinear effects are present.
+As shown in Figure 4 (left), NIMO perfectly recovers the ground truth coefficients, confirming that the neural network does not distort the linear coefficients, maintaining the intelligibility of MEM.
+We then extend the analysis to three additional regression settings (Setting 1, 2, 3) with varying nonlinearities and sparsity levels (see Appendix D.2 for details).
+Unlike Lasso, NIMO not only recovers the true coefficients but also correctly identifies the sparsity pattern, i.e., which features are uninformative (Figure 4, bottom-right).
+Based on these results, we conclude that NIMO estimates the accurate marginal effects at the mean and preserves the additive population-level contribution of each feature.
+Hypothesis 3:
+NIMO achieves a favorable trade-off between interpretability and predictive performance in practice.
+Figure 5: Predictive performance (MSE) the diabetes, Boston housing and superconductivity dataset.
+This experiment aims to evaluate whether NIMO can achieve strong predictive performance on real datasets and how the results can be interpreted.
+We select several widely used datasets from the UCI Machine Learning Repository: the diabetes dataset (Efron et al., 2004), the Boston housing dataset (Belsley et al., 2005), and the superconductivity dataset (Hamidieh, 2018).
+Each dataset is randomly shuffled and split multiple times with different seeds (9 seeds for the diabetes dataset and the Boston housing dataset, 5 seeds for the superconductivity dataset), and the experiments are performed once on each split using identical model settings.
+(a) Diabetes
+(b) Boston
+Figure 6: MEMs for Lasso and NIMO on the diabetes (left) and Boston housing dataset (right).
+We first examine the predictive performance of NIMO.
+The box plot of the mean squared error for each method is shown in Figure 5.
+NIMO generally achieves performance compatible with the other methods and attains the best performance on the superconductivity dataset.
+To assess interpretability, we first visualize the marginal effects at the mean (MEM) obtained from Lasso and NIMO for the diabetes and Boston housing datasets in Figure 6.
+Recall that MEM estimates marginal effects at the average values of the input features, reflecting the additive population-level contribution of each feature.
+The visualizations reveal that the MEM from NIMO differ substantially from those of Lasso, reflecting NIMO’s ability to capture nonlinear feature interactions that Lasso cannot represent.
+For the superconductivity dataset, which involves a much larger set of features, the corresponding MEM visualization is provided in the Appendix.
+We then use MEM to rank the features on the diabetes and Boston housing datasets.
+For comparison, we also apply the post-hoc explainer SHAP (Lundberg & Lee, 2017) to our model and rank features according to the global SHAP values.
+The results are shown in Table 2.
+On the diabetes dataset, all methods behave similarly in identifying the most important features, which is expected since the data lacks complex nonlinear interactions.
+This is further supported by the comparable predictive performance across methods, as shown in Figure 5 (left).
+Notably, SHAP produces identical feature rankings to MEM for NIMO.
+In contrast, for the Boston housing dataset, the methods yield inconsistent feature rankings, suggesting that nonlinear interactions play a crucial role in explaining the data.
+Although MEM can also be computed for other nonlinear methods, they do not provide the same level of intelligible interpretability as our model.
+For NIMO, SHAP rankings are largely consistent with MEM, though not perfectly aligned.
+The most pronounced difference appears with LassoNet, which assigns high importance to features considered least relevant by other methods, such as INDUS.
+We attribute this to how LassoNet handles nonlinearities.
+Table 2: Feature rankings according to MEMs on the diabetes and Boston housing datasets.
+A lower ranking is associated with a higher feature importance.
+Features whose rankings are consistently similar across multiple methods are highlighted in yellow.
+Diabetes Dataset
+Feature SHAP NIMO LassoNet NAM IMN Lasso age 7 7 9 5 9 10 sex 5 5 5 7 5 5 bmi 1 1 2 2 2 2 bp 3 3 3 3 3 4 s1 10 10 7 10 8 3 s2 9 9 10 9 7 6 s3 4 4 4 6 6 8 s4 8 8 8 8 4 7 s5 2 2 1 1 1 1 s6 6 6 6 4 10 9
+Boston Housing Dataset
+Feature SHAP NIMO LassoNet NAM IMN Lasso CRIM 10 9 3 5 8 9 ZN 13 13 10 11 12 8 INDUS 11 12 2 12 13 13 CHAS 12 8 5 8 10 11 NOX 8 11 9 10 9 7 RM 2 2 8 6 1 3 AGE 6 7 6 9 5 12 DIS 3 3 1 3 3 2 RAD 5 1 12 13 7 4 TAX 4 5 11 7 6 6 PTRATIO 7 6 4 4 4 5 B 9 10 13 2 11 10 LSTAT 1 4 7 1 2 1
+4.3 Ablation study
+To verify the effectiveness of masking out the $j$-th component of an instance when passing it through the neural network, we conducted an ablation study.
+Specifically, we examined the effect of masking out the $j$-th component as well as the effect of removing position encoding.
+From the ablation results across the three settings (Figure 7, Figure 15, and Figure 16), we observe several consistent patterns:
+  • Without masking the $j$-th component, the first fc layer of the neural network can still exhibit sparsity patterns, but they are less distinct than when masking is applied.
+We attribute this to the position encoding, which enables the network to identify the feature index.
+  • Without masking the $j$-th component, the MEMs are confounded by nonlinear interactions, so their interpretation as additive regression coefficients and the meaning of their magnitudes are lost.
+  • Further removing the position encoding distorts the sparsity patterns in the first fc layer and alters the magnitude of the coefficients, undermining interpretability.
+Figure 7: The effect of feature masking and position encoding on the linear coefficients and the sparsity of the first fc layer on setting 1.
+5 Conclusions
+In this paper, we propose a nonlinear interpretable model (NIMO).
+Unlike other hybrid approaches, our model preserves the interpretability of linear coefficients, capturing the additive contribution of each feature to the prediction, while also accommodating nonlinear interactions.
+Interpretability is formalized through marginal effects, with the marginal effects at the mean (MEM) providing global, population-level summaries that align directly with the linear coefficients.
+To enable effective training, we introduce an optimization algorithm based on parameter elimination, and enforce sparsity through adaptive ridge regularization.
+We demonstrate the effectiveness of the proposed approach on both synthetic and real datasets, showing competitive predictive performance while delivering meaningful and globally consistent interpretations.
+A key limitation of the approach is that it assumes the data can be described as a linear model with nonlinearity corrections; highly nonlinear datasets or features with self-interactions (e.g., $x_{j}f(x_{j})$) may not be fully captured.
+In such cases, careful feature engineering may be required to preserve interpretability.
+Appendix
+Appendix A Equivalence of Adaptive Ridge regression and Lasso
+The equivalence in the solution of Adaptive ridge regression and Lasso regression is a well-known fact (Tibshirani, 1996; Grandvalet, 1998; Grandvalet & Canu, 1998).
+Since both our model relies on this fact both in the regression and classification settings, we report its proof below.
+Proof.
+Let $X\in\mathbb{R}^{n\times d}$ be $n$ $d$-dimensional observations and $\mathbf{y}\in\mathbb{R}^{n}$ be the targets.
+In a linear model the relationship between target and data is assumed to be linear:
+$\mathbf{y}=X\bm{\beta}+\epsilon\;,$ (9)
+where $\bm{\beta}\in\mathbb{R}^{d}$ is the coefficient vector and $\epsilon\sim\mathcal{N}(0,1)$ is standard Gaussian noise. In Ridge Regression a penalized version of the objective is minimized:
+$\min_{\bm{\beta}}\|X\bm{\beta}-\mathbf{y}\|_{2}^{2}+\lambda\sum_{j=1}^{d}\beta_{j}^{2}\;,$ (10)
+where $\lambda$ weighs the penalty term.
+In Adaptive Ridge regression, instead of a single $\lambda$, a set of feature-specific penalties $\{\nu_{i}\}_{i=1}^{d}$ are introduced.
+The objective then reads as
+$\min_{\bm{\beta}\,\nu}\|X\bm{\beta}-\mathbf{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}\quad\text{s. t.}\quad\sum_{i=1}^{d}\frac{1}{\nu_{i}}=\frac{d}{\lambda}\;,$ (11)
+where $\nu_{j}\geq 0$ $\forall j$ and $\lambda$ is a predefined value. One can solve the above constrained optimization through the method of Lagrange multipliers:
+$\mathcal{L}(\bm{\nu},\mu)=\|X\bm{\beta}-\mathbf{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}+\mu\sum_{i=1}^{d}\frac{1}{\nu_{i}}\;,$ (12)
+where $\mu$ is the Lagrangian multiplier and $\bm{\nu}\coloneqq[\nu_{1},\ldots,\nu_{d}]^{T}$. The stationary points can be found by taking the gradients with respect to $\nu_{k}$, as follows:
+$\frac{\partial\mathcal{L}(\bm{\nu},\mu)}{\partial\nu_{k}}=\beta_{k}^{2}-\mu\nu_{k}^{-2}=0\quad\Rightarrow\quad\nu_{k}\beta_{k}^{2}=\mu/\nu_{k}\quad\text{or}\quad\nu_{k}|\beta_{k}|=\mu^{1/2}$ (13)
+If we now substitute $\nu_{k}|\beta_{k}|=\mu^{1/2}$ in the Adaptive Ridge objective in Eq. 11, we get the Lasso objective:
+$f(\bm{\beta},\bm{\nu})$ $=\|X\bm{\beta}-\mathbf{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}$ (14)
+$=\|X\bm{\beta}-\mathbf{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}|\beta_{i}|\cdot|\beta_{i}|$
+$=\|X\bm{\beta}-\mathbf{y}\|^{2}+\sum_{i=1}^{d}\mu^{1/2}|\beta_{i}|$
+$=\|X\bm{\beta}-\mathbf{y}\|^{2}+\mu^{1/2}\sum_{i=1}^{d}|\beta_{i}|.$
+which gives the relationship between adaptive ridge regression and Lasso regression, with the Lasso penalty parameter at optimum to be $\mu^{1/2}=\frac{\lambda}{d}\sum_{k}|\beta_{k}|$:
+[…]
+[…]
+$\Rightarrow$ $\quad\mu^{1/2}=\frac{\lambda}{d}\sum_{k}|\beta_{k}|$
+Appendix B Training via parameter elimination
+The proposed model in Eq.
+2 has two sets of parameters: the linear coefficients $\bm{\beta}$ and the neural network parameters $\bm{u}$.
+Since these two sets of parameters are tightly entangled, jointly optimizing them via gradient descent is not straightforward in practice.
+Furthermore, to increase the interpretability of the model, we also aim to achieve sparsity in the learned coefficients.
+In this section, we show how to achieve these optimization goals via parameter elimination.
+Specifically, we eliminate $\bm{\beta}$ by expressing it in terms of the neural network parameters $\bm{u}$ and substituting it back into the optimization objective.
+The resulting problem can then be optimized effectively and efficiently using gradient descent.
+Sparsity is achieved by reformulating the problem as an adaptive ridge regression.
+We further show that this approach can be extended to generalized linear models, and we illustrate it in detail for logistic regression.
+B.1 Regression setting: adaptive ridge regression
+We already showed that the proposed model in Eq. 2 can be re-written in matrix form as:
+$f(X)=B_{\bm{u}}\bm{\beta}\quad\text{with}\quad B_{\bm{u}}=X+X\circ G_{\bm{u}}$ (16)
+where $\bm{\beta}\in\mathbb{R}^{d}$ are the linear coefficients, $G_{\bm{u}}\in\mathbb{R}^{n\times d}$ is the neural network output with inputs $X\in\mathbb{R}^{n\times d}$ and $\circ$” denotes element-wise multiplication.
+The linear model part is then expressed as:
+$\mathbf{y}=B_{\bm{u}}\bm{\beta}+\epsilon,$ (17)
+where $\mathbf{y}\in\mathbb{R}^{n}$ denotes the targets. Adaptive Ridge regression minimizes the following objective:
+$\min_{\bm{\beta},\bm{u}}\|B_{\bm{u}}\bm{\beta}-\mathbf{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}\quad\text{s. t.}\quad\sum_{i=1}^{d}\frac{1}{\nu_{i}}=\frac{d}{\lambda}$ (18)
+where $\nu_{i}>0$ are the penalty parameters for each coefficient $\beta_{i}$, and $\lambda>0$ is a predefined parameter.
+To avoid divergent solutions we use the re-parameterization employed in Grandvalet (1998); Grandvalet & Canu (1998):
+$\gamma_{i}=(\nu_{i}/\lambda)^{1/2}\beta_{i},\quad c_{i}=(\lambda/\nu_{i})^{1/2}\;.$ (19)
+If we substitute $\beta_{i}$ and $\nu_{i}$ into the Adaptive Ridge regression objective, we get the following optimization problem:
+$\min_{\bm{\gamma},\bm{c},\bm{u}}\|\mathbf{y}-B_{\bm{u}}D_{\bm{c}}\bm{\gamma}\|^{2}+\lambda\|\bm{\gamma}\|^{2}\quad\text{s. t.}\quad\sum_{i=1}^{d}c_{i}^{2}=d,\quad c_{i}\geq 0,$ (20)
+where $D_{\bm{c}}$ is a diagonal matrix whose diagonal elements are given by the vector $\bm{c}\coloneqq[c_{1},\ldots,c_{d}]^{T}$.
+Applying the parameter elimination trick, if we fix $\bm{c}$ and $\bm{u}$, $\bm{\gamma}$ can be expressed in a closed-form in terms of $\bm{c}$ and $\bm{u}$:
+$\hat{\bm{\gamma}}(\bm{c},\bm{u})=\left(D_{\bm{c}}B_{\bm{u}}^{T}B_{\bm{u}}D_{\bm{c}}+\lambda I\right)^{-1}D_{\bm{c}}B_{\bm{u}}^{T}\mathbf{y},$ (21)
+We can now substitute $\hat{\bm{\gamma}}(\bm{c},\bm{u})$ back into the original problem and take the constraint into consideration.
+This leads us to a Lagrangian minimization problem in the scaling variable $\bm{c}$ and neural network parameters $\bm{u}$:
+$\min_{\bm{c},\bm{u}}\|\mathbf{y}-B_{\bm{u}}D_{\bm{c}}\hat{\bm{\gamma}}(\bm{c},\bm{u})\|^{2}+\lambda\|\hat{\bm{\gamma}}(\bm{c},\bm{u})\|^{2}+\mu\|\bm{c}\|^{2},$ (22)
+where $\mu$ is the Lagrangian multiplier introduced for $\bm{c}$, and we can optimize $\bm{c},\bm{u}$ together through gradient descent.
+The training algorithm for this re-parameterized adaptive ridge regression is summarized in Alg.
+1.
+Algorithm 1 Training of Re-parameterized Adaptive Ridge Regression
+0: Training data $X\in\mathbb{R}^{n\times d}$, training target $\mathbf{y}\in\mathbb{R}^{n}$, neural network $g_{\bm{u}}(\cdot)$, regularization parameter $\lambda$ and $\mu$, learning rate $\eta$, maximum iterations $T$.
+1: Initialize network parameters $\bm{u}$ and scaling coefficients $\bm{c}$ (with $c_{i}\geq 0$).
+2:for $t=1,\dots,T$ do
+3:  $B_{\bm{u}}=X+X\circ G_{\bm{u}}$ // Compute modified design matrix
+4:  $D_{\bm{c}}=\operatorname{diag}(\bm{c})$ // Form diagonal scaling matrix
+5:  $\hat{\bm{\gamma}}=\bigl(D_{\bm{c}}B_{\bm{u}}^{T}B_{\bm{u}}D_{\bm{c}}+\lambda I\bigr)^{-1}D_{\bm{c}}B_{\bm{u}}^{T}\mathbf{y}$ // Solve for $\hat{\bm{\gamma}}$
+6:  $\hat{\bm{\beta}}=\bm{c}\circ\hat{\bm{\gamma}}$ // Compute regression coefficients
+7:  $L=\|\mathbf{y}-B_{\bm{u}}\hat{\bm{\beta}}\|^{2}+\lambda\|\hat{\bm{\gamma}}\|+\mu\|\bm{c}\|^{2}$ // Compute loss function
+8:  $\bm{c}\leftarrow\bm{c}-\eta\nabla_{\bm{c}}L,\quad\bm{u}\leftarrow\bm{u}-\eta\nabla_{\bm{u}}L$ // Gradient descent updates
+9:end for
+10:Return $\hat{\bm{\beta}},\bm{c},\bm{u}$ // Output final parameters
+B.2 Logistic regression: Iteratively reweighted least squares
+The procedure followed in the previous section can be extended to generalized linear models.
+Here we showcase how to do it for logistic regression.
+Logistic regression is a classification model that describes the probability that a binary outcome $y\in\{0,1\}$ occurs given a feature vector $\mathbf{x}\in\mathbb{R}^{d}$:
+$P(y=1|\mathbf{x})=\sigma(\mathbf{x}^{T}\bm{\beta})$ (23)
+where $\sigma(z)=\frac{1}{1+e^{-z}}$ is the sigmoid function.
+The model can then be trained by minimizing the negative log-likelihood.
+Similarly to the regression case discussed in the previous section, we add an adaptive ridge regression penalty to the objective:
+$\mathcal{L}(\bm{\beta})=-\sum_{i=1}^{n}[y_{i}\log\sigma(\mathbf{x}_{i}^{T}\bm{\beta})+(1-y_{i})\log(1-\sigma(\mathbf{x}_{i}^{T}\bm{\beta}))]+\sum_{j=1}^{d}\nu_{j}\beta_{j}^{2}\quad\text{ s.
+t.
+}\quad\sum_{j=1}^{d}\frac{1}{\nu_{j}}=\frac{d}{\lambda}$ (24)
+Now we can use the same re-parameterization used in the previous section for regression. We get the following standard ridge-penalized logistic regression:
+$\begin{gathered}\min_{\bm{\gamma},\bm{c}}-\sum_{i=1}^{n}\left[y_{i}\log\sigma\left((\mathbf{x}_{i}\odot\bm{c})^{T}\bm{\gamma}\right)+(1-y_{i})\log\left(1-\sigma\left((\mathbf{x}_{i}\odot\bm{c})^{T}\bm{\gamma}\right)\right)\right]+\lambda\|\bm{\gamma}\|^{2}\\ \text{ s.t.
+}\quad\sum_{j=1}^{d}c_{j}^{2}=d,\quad c_{j}\geq 0,\end{gathered}$ (25)
+where $\bm{c}:=[c_{1},...,c_{d}]^{T}$, $\odot$ is element-wise multiplication.
+The same as what we done for regression, we can eliminate $\bm{\gamma}$ by expressing it in a closed-form with respect to $\bm{c}$.
+To get this closed-form expression of $\bm{\gamma}$, we use the iteratively reweighted least squares (IRLS) algorithm.
+The associated update reads as:
+$\bm{\gamma}^{(t+1)}(\bm{c})=(\tilde{X}^{T}W^{(t)}\tilde{X}+\lambda I)^{-1}\tilde{X}^{T}W^{(t)}\mathbf{z}^{(t)}$ (26)
+where $\tilde{X}$ is the modified design matrix with entries $\tilde{x}_{ij}=c_{j}x_{ij}$, $W^{(t)}$ is a diagonal weight matrix with entries $w_{i}^{(t)}=p_{i}^{(t)}(1-p_{i}^{(t)}),p_{i}^{(t)}=\frac{1}{1+e^{-\tilde{\bm{x}}_{i}^{T}\bm{\gamma}^{(t)}}}$, and $\mathbf{z}^{(t)}$ is the the vector of working response with entries $z_{i}^{(t)}=\tilde{\bm{x}}_{i}^{T}\bm{\gamma}^{(t)}+\frac{y_{i}-p_{i}^{(t)}}{w_{i}^{(t)}}$.
+We then substitute $\bm{\gamma}^{(t+1)}(\bm{c})$ back into the original problem, resulting the optimization over $\bm{c}$ only.
+Note that in our model, our design matrix $X$ contains the nonlinear term as well, so we would need to replace $X$ with $B_{\bm{u}}=X+X\circ G_{\bm{u}}$.
+The rest goes the same as the case for regression.
+B.3 Extension to sub-$\ell_{1}$ pseudo-norms
+In some cases sparsity is a crucial aspect in the learning process.
+One of the limitations of lasso is the well-known over-shrinkage effect, which means that shrinkage of uninformative features could be achieved at the expense of shrinking the retained features as well.
+One way to avoid this is to use sub-$\ell_{1}$ pseudo-norms, which in the limits of the $\ell_{0}$ pseudo-norm achieve sparsity without any shrinkage.
+The main challenge with sub-$\ell_{1}$ pseudo-norm is that the resulting objective function becomes non-convex and harder to optimize.
+However, recent work has successfully used sub-$\ell_{1}$ norms as a prior on regression problems (Negri et al., 2023), which is similar to our setting.
+We now show how to use a sub-$\ell_{1}$ prior in our Adaptive Ridge regression objective, which we report below for completeness:
+$\min_{\bm{\beta}}\|X\bm{\beta}-\mathbf{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}\quad\text{s.t.}\quad\sum_{i=1}^{d}\frac{1}{\nu_{i}}=\frac{d}{\lambda}$ (27)
+In order to generalize to sub-$\ell_{1}$ pseudo-norms, we need the following simple modification on $\nu_{i}$:
+$\sum_{i=1}^{d}\frac{1}{\nu_{i}^{\delta}}=C,\quad 0<\delta\leq 1.$ (28)
+We now apply the same re-parameterization:
+$\gamma_{i}=(\nu_{i}/\lambda)^{1/2}\beta_{i},\quad c_{i}=(\lambda/\nu_{i})^{1/2}\;.$ (29)
+The objective function then reads as
+$\min_{\bm{\gamma},\bm{c}}\|\mathbf{y}-XD_{\bm{c}}\bm{\gamma}\|^{2}+\lambda\|\bm{\gamma}\|^{2}\quad\text{s.t.}\quad\sum_{i=1}^{d}c_{i}^{2\delta}=\lambda^{\delta}C,\quad c_{i}>0\;.$ (30)
+Fixing $\bm{c}$, we can eliminate $\bm{\gamma}$ by expressing it in a closed-form in terms of $\bm{c}$:
+$\hat{\bm{\gamma}}(\bm{c})=\left(D_{\bm{c}}X^{T}XD_{\bm{c}}+\lambda I\right)^{-1}D_{\bm{c}}X^{T}\mathbf{y}.$ (31)
+Substituting it back into the original optimization problem and taking the constraint into consideration lead us to a Lagrangian minimization problem in the scaling variable $\bm{c}$ only:
+$\min_{\bm{c}}\|\mathbf{y}-XD_{\bm{c}}\hat{\bm{\gamma}}(\bm{c})\|^{2}+\lambda\|\hat{\bm{\gamma}}(\bm{c})\|+\mu\sum_{i=1}^{d}c_{i}^{2\delta},$ (32)
+and we can optimize it through gradient descent.
+To verify that this constraint leads sub-$\ell_{1}$ pseudo-norm to the original $\bm{\beta}$, first we reformulate the original objective with the method of Lagrangian multipliers:
+$\mathcal{L}(\bm{\nu},\mu)=\|X\bm{\beta}-\mathbf{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}+\mu\sum_{i=1}^{d}\frac{1}{\nu_{i}^{\delta}}$ (33)
+Then, we find stationary points for $\nu_{i}$ by taking the gradients and setting them to zero:
+$\frac{\partial\mathcal{L}(\bm{\nu},\mu)}{\partial\nu_{i}}=\beta_{i}^{2}-\mu\delta\nu_{i}^{-(\delta+1)}=0\quad\Rightarrow$ $\quad\nu_{i}^{\delta+1}=\frac{\mu\delta}{\beta_{i}^{2}}$ (34)
+$\Rightarrow$ $\quad\nu_{i}=\left(\frac{\mu\delta}{\beta_{i}^{2}}\right)^{\frac{1}{\delta+1}}$
+The penalty in the original adaptive ridge regression can be rewritten as:
+$\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}=\sum_{i=1}^{d}\left(\frac{\mu\delta}{\beta_{i}^{2}}\right)^{\frac{1}{\delta+1}}\beta_{i}^{2}\propto\sum_{i=1}\beta_{i}^{2(1-\frac{1}{\delta+1})}.$ (35)
+Note that when $\delta=1$, we recover the Lasso $\ell_{1}$ norm. When $\delta\rightarrow 0$, it approaches the $\ell_{0}$ pseudo-norm, which basically counts the number of nonzero entries.
+Appendix C NIMO: Implementation details
+In this section, we provide a detailed introduction to the structure of NIMO, along with the strategies we used to improve training.
+C.1 Model Definition
+We define NIMO as follow:
+$f(\mathbf{x})=\beta_{0}+\sum_{j=1}^{d}x_{j}\beta_{j}(1+g_{\bm{u}_{j}}(\mathbf{x}_{-j}))\quad\text{s.t.}\quad g_{\bm{u}_{j}}(\mathbf{0})=0\;.$ (36)
+The architecture of our model consists of two components, as illustrated in Fig.
+2.
+The left component is a nonlinear neural network that operates on the input features $\mathbf{x}_{-j}$, where the $j$-th element has been masked out.
+The right component is a linear model, in which coefficients $\beta_{j}$ are multiplied by the corresponding neural network outputs and subsequently summed.
+In practice, for each feature component $x_{j}$ of the input vector $\mathbf{x}$, the value is temporarily masked to form $\mathbf{x}_{-j}$, which is then passed through a neural network $g_{\bm{u}_{j}}$.
+The output of this network is added to a constant $1$ and multiplied by the original feature value $x_{j}$, resulting in a rescaled version of the feature:
+$\tilde{x}_{j}=x_{j}\cdot(1+g_{\bm{u}_{j}}(\mathbf{x}_{-j}))$ (37)
+This rescaling can be interpreted as an element-wise modulation of the input features based on context.
+The rescaled features are then linearly combined using fixed coefficients $\bm{\beta}$ to produce the final prediction:
+$y=\beta_{0}+\sum_{j=1}^{d}\tilde{x}_{j}\beta_{j}.$
+C.2 Practical Implementation
+We implement our model with PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019).
+While the model definition is simple, in practice, several strategies are needed to make the training more efficient and effective.
+Positional Encoding
+In principle each feature $x_{j}$ could have its own dedicated network $g_{\bm{u}_{j}}$, but this can be computationally expensive for high-dimensional inputs.
+To reduce overhead, we instead use a shared neural network $g_{\bm{u}}$ across all components.
+To ensure the network remains aware of which feature is currently masked out, we append a positional encoding to $\mathbf{x}_{-j}$ before passing it through the shared network.
+We assign each masked input feature $\mathbf{x}_{-j}$ a unique binary vector corresponding to its index $j$, represented in fixed-length binary form using $\lfloor\log_{2}d\rfloor+1$ bits.
+This compact encoding allows the shared neural network to remain index-aware with minimal overhead.
+Group Penalty and Noise Injection
+In the main paper we showed that linear coefficients $\bm{\beta}$ in NIMO are interpretable as in a linear model.
 Even though the neural network contribution is not interpretable per se, we can still identify relevant nonlinear interactions by means of sparsity.
 The idea is to force the first layer of the neural network to be sparse, hence to use only relevant features.
-To do so, we apply a group penalty on the weight matrix $W\in\mathbb{R}^{p\times d}$ of the first fully-connected (fc) layer, where $d$ is the number of input features and $p$ the hidden dimension of the first fc layer.
-In order to obtain sparsity at a input feature level, we need to consider neurons connected to such feature and penalize them as a group:
-$\ell_{\text{group}}=\lambda_{\text{group}}\sum_{j=1}^{d}\|\mathbf{w}_{j}\|_{2}\;.$ (6)
+To do so, we apply a group penalty on the weight matrix $W\in\mathbb{R}^{p\times d}$ of the first fully-connected (fc) layer, where $d$ is the number of input features and $p$ is the hidden dimension of the first fc layer.
+In order to obtain sparsity at the input feature level, we need to consider neurons connected to the same input feature and penalize them as a group:
+$\ell_{\text{group}}=\lambda_{\text{group}}\sum_{j=1}^{d}\|\mathbf{w}_{j}\|_{2}\;,$ (38)
 where $\mathbf{w}_{j}=W[:,j]\in\mathbb{R}^{p}$ is the weight vector acting on the $j$-th feature.
@@
 By doing so, we disrupt the deterministic signal path, making it harder for the subsequent layers to simply exploit the resulting weak, regularized, and noisy signal.
-As a result, the network is encouraged to focus on strong and robust signals associated with informative input features.This approach works very well in practice and allows to easily achieve sparsity.
-We provide further details in the Appendix.
-3.4 Limitations
-The proposed approach is designed to maintain the interpretability of linear models and, at the same time, to learn nonlinear corrections to linear predictions.
-Therefore, the underlying assumption is that the data can be described by adding nonlinearities over a linear model.
-Highly nonlinear datasets might not be suitable for the proposed approach.
-However, in such cases the question itself about interpretability would be ill-posed, at least according to our notion of interpretability.
-There is however one case that cannot be modeled by Eq.
-(2).
-Assume the underlying data generating process contains a self-interacting term of the kind $x_{j}f(x_{j})$, where $f$ is a linear or nonlinear function (e.g.
-$x\sin(x)$ or $x^{2}$).
-In order to make the linear coefficients interpretable we assumed $g_{\bm{u}_{j}}(\bm{\bm{x}_{-j}})$ not to depend on $x_{j}$, which clearly prevents us from learning such types of nonlinearities.
-However, if such feature is known to play a role in the dataset, simple feature engineering can solve the issue.
-4 Experiments
-In this section, we experimentally showcase the proposed approach.
-Firstly, in Section 4.1 we use a toy example to illustrate how the model works.
-Secondly, in Section 4.2 we show that across several synthetic datasets the model learns the actual underlying coefficients, the nonlinearities and the correct sparsity in the networks.
-We also achieve higher performance compared to other hybrid approaches.
-Lastly, in Section 4.3 we test the model on real datasets and show that it is able to select a sparser set of features than Lasso, while maintaining consistency with Lasso in terms of the signs and magnitudes of the selected coefficients.
-However, thanks to the nonlinear terms the model achieves significantly higher performance, on par with neural networks but with the benefits of inherent interpretability.
-In the Appendix we provide further details on each experimental setup.
-Figure 3: Illustrative toy example.
-Left: NIMO learns the GT coefficients and also finds the correct sparsity.
-Middle: the learned nonlinearities coincide with the GT.
-Right: the first fc layer of the NN learns the correct sparsity: only features $x_{1}$ and $x_{2}$ have non-zero weights, while $x_{3}$ is ignored.
-4.1 Toy example
-We first illustrate NIMO on a simple three-dimensional toy example. We generate the data as follows:
-$y=3\cdot x_{1}\cdot(1+\tanh(10x_{2}))+(-3)\cdot x_{2}\cdot(1+\sin(-2x_{1}))+0\cdot x_{3}+\epsilon\;,$ (7)
-where $\epsilon$ is the usual Gaussian noise.
-This toy example has 3 features $\{x_{`},x_{2},x_{3}\}$, but $x_{3}$ is actually inactive (hence uninformative) since it is multiplied by a zero coefficient.
-We generate 400 samples and use a 200-100-100 train-validation-test split.
-As a baseline, we compare with Lasso regression.
-As shown in Figure 3 (left) the $\bm{\beta}$ coefficients learned with NIMO coincide with the ground truth.
-Relevantly, NIMO correctly learns $\beta_{3}$ to be zero, i.e.
-that $x_{3}$ is uninformative.
-In Figure 3 (middle) we show that the neural network used in NIMO can perfectly recover the underlying nonlinear interactions both for $x_{1}$ and $x_{2}$.
-Furthermore, if we analyze the weights of the first fc layer, we can see that the neurons connected to the input $x_{3}$ are sparse; see Figure 3 (right).
-This provides further interpretability in the model because it implies that $x_{3}$ does not contribute to nonlinear interactions.
-4.2 Synthetic Experiments
-We now showcase that NIMO is able to recover sparse coefficients and learn the underlying nonlinearities across a set of different regression and classification settings, with different number of features and sparsity levels.
-We compare out method against Lasso regression (Tibshirani, 1996), a vanilla neural network and two hybrid approaches, namely LassoNet (Lemhadri et al., 2021), Contextual Lasso (Thompson et al., 2023).
-Contextual Lasso requires the features to be separated in explanatory and contextual, which is not directly applicable to our settings since we do not have contextual features.
-To still compare with Contextual Lasso, we use all features as context.
-Compared with other hybrid approaches, NIMO achieves better or equivalent performance while still retaining interpretability.
-In all settings we use 200 training samples and 100 for validation and test.
-Further details can be found in the Appendix.
-Figure 4: Learned $\bm{\beta}$ coefficients on synthetic regression datasets and comparison with Lasso regression and Ground truth.
-Vanilla regression (left) is purely linear while Settings 1, 2, 3 (right) have different nonlinearities.
-Both NIMO and Lasso recover the correct nonzero coefficients (informative features) but only NIMO recovers the correct sparsity of the zero coefficients (uninformative features).
-Regression We design synthetic regression datasets by explicitly controlling the linear coefficients and the nonlinearities, similarly to Eq.
-(7).
-We explore multiple nonlinearities, different dimensionality and also use several uninformative features (i.e.
-with zero coefficient) to test if we can recover sparse signals.
-As a sanity check, we first test NIMO on a Vanilla regression dataset with linear coefficients in $\{-5,-4,...,3,4\}$ but without nonlinearities.
-The idea is to check if the neural network of NIMO interferes with the linear part in the absence of nonlinearities.
-As shown in Figure 4 (left), NIMO is able to perfectly recover the ground truth coefficients, and the neural network does not interfere with them.
-We provide three more regression settings (Setting 1, 2, 3) where we test different nonlinearities and different sparsity levels.
-We provide details about the exact data generating process in the Appendix.
-In Table 1 we report the MSE on the test set over all settings for NIMO and we compare it against linear, nonlinear and hybrid approaches.
-Overall, NIMO achieves the best performance across all settings with nonlinearities.
-This is possible because NIMO learns the correct coefficients (Figure 4) and, at the same time, accurately learns the underlying nonlinear interactions.
-Differently from Lasso, we can also perfectly learn the sparsity pattern in the coefficients, i.e.
-which features are uninformative; see Figure 4 (bottom, right).
-In contrast, nonlinear models and hybrid models learn spurious nonlinear relationships.
-Table 1: MSE loss for synthetic regression settings (see the Appendix for more details). We compare with Lasso regression, NN, LassoNet and Contextual Lasso. In each setting we use 200 samples.
-Features Lasso NN NIMO LassoNet Contextual Lasso
-Vanilla regr. 10 0.010 0.939 0.048 0.015 -
-Toy Example 3 18.982 0.446 0.166 0.628 0.646
-Setting 1 5 3.164 1.109 0.030 0.078 0.381
-Setting 2 10 3.340 1.482 0.217 2.991 3.418
-Setting 3 50 13.122 13.718 0.334 1.342 13.070
-Classification Our method can be extended to generalized linear models and here we showcase it for logistic regression.
-In order to create data for synthetic classification we follow a similar procedure as in the regression case, where we directly control the linear coefficients and the nonlinear interactions.
-The only difference is that we further apply a link function and binarize the output and get the labels.
-Note that this way we loose information about the original linear coefficients, which cannot be recovered exactly anymore.
-The sparsity in the coefficients still remains crucial.
-Also in this case we show various settings for different dimensionality, nonlinear terms and different number of uninformative features.
-We provide more details about each setting in the Appendix.
-In Table 2 we report the classification accuracy on the test set for all methods.
-Despite our efforts, we were not able to train Contextual lasso on such classification settings, even when all features were given as contextual features.
-Overall, our model significantly outperforms logistic regression on all settings.
-When analyzing the learned coefficient, which for brevity we report in the Appendix, we can see that only NIMO recovers the correct sparsity patterns.
-As argued before, this allows to learn the correct nonlinear interactions.
-Other nonlinear models and hybrid approaches perform similarly to NIMO but fail to provide the interpretability of the coefficients as in a linear model.
-Table 2: Classification accuracy for synthetic classification settings (see the Appendix for more details). We compare with logistic regression, NN and LassoNet. In each setting we use 200 samples.
-Features Log. Regr. NN NIMO LassoNet
-Setting 1 3 0.59 0.90 0.92 0.89
-Setting 2 10 0.68 0.83 0.85 0.82
-Setting 3 50 0.74 0.69 0.84 0.90
-Figure 5: Diabetes datasets results.
-Left: MSE loss on test set.
-All methods achieve comparable results.
-Middle: Learned linear coefficients by Lasso and NIMO.
-NIMO selects sparser coefficients (“s1, “s2”, “s4”).
-Right: the first fc layer of NIMO is very sparse; few nonlinearities are relevant.
-Errorbars are obtained by running all models across 5 different datasets splits.
-4.3 Real Datasets
-We further evaluated our model on two real-world datasets: the diabetes dataset (Efron et al., 2004) and the Boston housing dataset (Belsley et al., 2005).
-In all the experiments, we use a 60%-20%-20% train-validation-test split.
-We compare our model with Lasso regression, a standard neural network and LassoNet (Lemhadri et al., 2021) in terms of MSE loss and feature sparsity.
-Diabetes dataset
-The diabetes dataset is a classic regression dataset consisting of 442 patients, each with 10 numeric features, and the target measures disease progression one year after baseline.
-As shown in Figure 5 (left), we observe that all methods achieve similar MSE losses and the neural network (also the hybird methods) shows no clear advantage on this dataset.
-Since the diabetes dataset is commonly used for linear regression tasks, it is not surprising that the data does not exhibit complex nonlinear interactions among features.
-We can verify this claim by looking at the weight sparsity of the first fc layer in NIMO, shown in Figure 5 (right).
-Relevantly, NIMO achieves a significantly sparser solution than Lasso regression, see Figure 5 (middle).
-This is particularly evident by looking at the coefficients associated with the feature “s1”, “s2” and “s4”.
-Boston dataset
-The Boston housing dataset contains 506 instances, each with 13 features, and the target is the median value of owner-occupied homes.
-As shown in Figure 6, our method performs significantly better than Lasso regression and shows compatible MSE test loss with the other nonlinear and hybrid approaches.
-This suggests that nonlinear interactions are crucial to explain the data.
-While both the neural network and LassoNet are able to capture those nonlinearities, they both lack the interpretability provided by our model.
-It is important to note that the Boston housing dataset has been criticized for containing controversial features  [^1] https://fairlearn.org/main/user_guide/datasets/boston_housing_data.html.
-The feature of interest is termed “B”.
-From the coefficient plot in Figure 6 (middle), we observe that the significance of the feature “B” is greatly reduced in our model compared to Lasso regression.
-This surprising result suggests that the reason why such feature appeared to be relevant is due to some artifacts of linear models.
-Thanks to the decoupling between the linear coefficients and the nonlinearities in NIMO, our results suggests that the role of the feature “B” towards the prediction might be irrelevant.
-Note that the feature “B” is also not relevant for the nonlinear terms, as shown in Figure 6.
-Figure 6: Boston datasets results.
-Left: Comparison of MSE losses.
-NIMO performs on par with other methods and significantly outperforms Lasso, suggesting nonlinear interactions are crucial.
-Middle: Learned linear coefficients by Lasso and NIMO.
-The two models learn coefficient with a comparable sparsity.
-Relevantly, the feature “B” is almost irrelevant for NIMO.
-Right: sparsity in the first fc layer of NIMO.
-Errorbars are obtained by running all models over 5 datasets splits.
-5 Conclusions
-In this paper we propose a nonlinear interpretable model.
-Different from other hybrid approaches, our model retains the interpretability of linear coefficients – just as in standard linear models – while also allowing for nonlinear interaction terms.
-However, unlike linear models, we are not limited to global linear coefficients that operate at the population level.
-Our model can also learn local nonlinearities, enabling the generation of modified linear coefficients tailored to individual data points.
-This can be viewed as an individual correction on top of the linear prediction.
-In order to effectively train the proposed approach, we propose an optimization algorithm based on profile likelihood and adaptive ridge regression.
-Specifically, we are able to optimize the neural network parameters via gradient descent while analytically updating the linear coefficients.
-By means of sparsity in the neural networks we can also automatically select relevant nonlinear interactions.
-We showcase the proposed approach on synthetic dataset and show that we can learn the underlying coefficients and nonlinearities, outperforming nonlinear and hybrid approaches.
-We also showcase the model on real world datasets and confirmed the same findings.
-When the dataset does not contain nonlinear interaction terms we recover sparse coefficients and the network becomes less relevant.
-When the dataset contains nonlinearities we can learn them with the same accuracy of other nonlinear approaches, but with the significant advantage of learning the interpretable linear coefficients as well.
-Appendix
-[…]
-[…]
-Appendix A Equivalence of Adaptive Ridge regression and Lasso
-The equivalence in the solution of Adaptive ridge regression and Lasso regression is a well-known fact [Tibshirani, 1996, Grandvalet, 1998, Grandvalet and Canu, 1998].
-Since both our model relies on this fact both in the regression and classification settings, we report its proof below.
-Proof.
-Let $X\in\mathbb{R}^{n\times d}$ be $n$ $d$-dimensional observations and $\bm{y}\in\mathbb{R}^{n}$ be the targets. In a linear model the relationship between target and data is assumed to be linear:
-$\bm{y}=X\bm{\beta}+\epsilon\;,$ (8)
-where $\bm{\beta}\in\mathbb{R}^{d}$ is the coefficient vector and $\epsilon\sim\mathcal{N}(0,1)$ is standard Gaussian noise. In Ridge Regression a penalized version of the objective is minimized:
-$\min_{\bm{\beta}}\|X\bm{\beta}-\bm{y}\|_{2}^{2}+\lambda\sum_{j=1}^{d}\beta_{j}^{2}\;,$ (9)
-where $\lambda$ weighs the penalty term.
-In Adaptive Ridge regression, instead of a single $\lambda$, a set of feature-specific penalties $\{\nu_{i}\}_{i=1}^{d}$ are introduced.
-The objective then reads as
-$\min_{\bm{\beta}}\|X\bm{\beta}-\bm{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}\quad\text{s. t.}\quad\sum_{i=1}^{d}\frac{1}{\nu_{i}}=\frac{d}{\lambda}\;,$ (10)
-where $\nu_{j}\geq 0$ $\forall j$ and $\lambda$ is a predefined value. One can solve the above constrained optimization through the method of Lagrange multipliers:
-$\mathcal{L}(\bm{\nu},\mu)=\|X\bm{\beta}-\bm{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}+\mu\sum_{i=1}^{d}\frac{1}{\nu_{i}}\;,$ (11)
-where $\mu$ is the Lagrangian multiplier and $\bm{\nu}\coloneqq[\nu_{1},\ldots,\nu_{d}]^{T}$. The stationary points can be found by taking the gradients with respect to $\nu_{k}$, as follows:
-$\frac{\partial\mathcal{L}(\bm{\nu},\mu)}{\partial\nu_{k}}=\beta_{k}^{2}-\mu\nu_{k}^{-2}=0\quad\Rightarrow\quad\nu_{k}\beta_{k}^{2}=\mu/\nu_{k}\quad\text{or}\quad\nu_{k}|\beta_{k}|=\mu^{1/2}$ (12)
-We can now use the constraint in Eq. (10):
-[…]
-[…]
-$\Rightarrow$ $\quad\mu^{1/2}=\frac{\lambda}{d}\sum_{k}|\beta_{k}|$
-If we now substitute $\nu_{k}|\beta_{k}|=\mu^{1/2}$ in the Adaptive Ridge objective in Eq. (10), we get the Lasso objective:
-$f(\bm{\beta},\bm{\nu})$ $=\|X\bm{\beta}-\bm{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}$ (14)
-$=\|X\bm{\beta}-\bm{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}|\beta_{i}|\cdot|\beta_{i}|$
-$=\|X\bm{\beta}-\bm{y}\|^{2}+\sum_{i=1}^{d}\mu^{1/2}|\beta_{i}|$
-$=\|X\bm{\beta}-\bm{y}\|^{2}+\mu^{1/2}\sum_{i=1}^{d}|\beta_{i}|.$
-which gives the relationship with the lasso turns out to be a Lasso regression, with the Lasso penalty parameter at optimum to be $\mu^{1/2}=\frac{\lambda}{d}\sum_{k}|\beta_{k}|$. ∎
-Appendix B Training through profile likelihood
-The proposed model in Eq.
-(2) has two sets of parameters: the linear coefficients $\bm{\beta}$ and the neural network parameters $\bm{u}$.
-Since the neural network learn the nonlinear corrections over the linear predictions, the two sets of parameters are clearly entangled and in practice optimizing them both with gradient descent is tricky.
-Furthermore, to increase the interpretability of the model we also encourage sparse coefficients and a sparse first fc layer of the the neural network.
-In this section we show how to make the optimization more effective and, at the same time, how to encourage sparsity.
-In particular, we propose a novel optimization method based on profile likelihood with adaptive ridge regression.
-The profile likelihood allows for an analytical update of the linear coefficients while adaptive rigde regression provides sparsity as in Lasso regression.
-We further show that we can extend such an approach to generalized linear models and we showcase it in detail for logistic regression.
-B.1 Regression setting: adaptive ridge regression
-Let $\mathcal{L}(\bm{\beta},\bm{u})$ be the joint likelihood in Eq.
-(3).
-First, we fix the neural network parameters and analytically solve the maximization problem to obtain $\hat{\bm{\beta}}$.
-Then, we substitute $\hat{\bm{\beta}}$ back into the joint likelihood, which results in the profile likelihood for $\bm{u}$: $\mathcal{L}(\hat{\bm{\beta}},\bm{u})$.
-We can then optimize $\bm{u}$ by minimizing the negative log-profile likelihood using gradient descent.
-We iterate over these steps until convergence.
-More specifically, we draw the inspiration from the Adaptive Ridge (AR) regression and exploit its equivalence to Lasso regression [Tibshirani, 1996, Grandvalet, 1998, Grandvalet and Canu, 1998].
-We already showed that the proposed model in Eq. (2) can be re-written in matrix form as:
-$f(X)=B_{\bm{u}}\bm{\beta}\quad\text{with}\quad B_{\bm{u}}=X+X\circ G_{\bm{u}}$ (15)
-where $\bm{\beta}\in\mathbb{R}^{d}$ are the linear coefficients, $G_{\bm{u}}\in\mathbb{R}^{n\times d}$ is the neural network output with inputs $X\in\mathbb{R}^{n\times d}$ and $\circ$” denotes element-wise multiplication.
-The linear model part is then expressed as:
-$\bm{y}=B_{\bm{u}}\bm{\beta}+\epsilon,$ (16)
-where $\bm{y}\in\mathbb{R}^{n}$ denotes the targets. Adaptive Ridge regression minimizes the following objective:
-$\min_{\bm{\beta}}\|B_{\bm{u}}\bm{\beta}-\bm{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}\quad\text{s. t.}\quad\sum_{i=1}^{d}\frac{1}{\nu_{i}}=\frac{d}{\lambda}$ (17)
-where $\nu_{j}>0$ are the penalty parameters for each coefficient $\beta_{j}$, and $\lambda>0$ is a predefined parameter.
-To avoid divergent solutions we use the re-parameterization employed in Grandvalet [1998], Grandvalet and Canu [1998]:
-$\gamma_{i}=(\nu_{i}/\lambda)^{1/2}\beta_{i},\quad c_{i}=(\lambda/\nu_{i})^{1/2}\;.$ (18)
-If we substitute $\beta_{i}$ and $\nu_{i}$ into the Adaptive Ridge regression objective, we get the following optimization problem:
-$\min_{\bm{\gamma}}\|\bm{y}-B_{\bm{u}}D_{\bm{c}}\bm{\gamma}\|^{2}+\lambda\|\bm{\gamma}\|^{2}\quad\text{s. t.}\quad\sum_{i=1}^{d}c_{i}^{2}=d,\quad c_{i}>0,$ (19)
-where $D_{\bm{c}}$ is a diagonal matrix whose diagonal elements are given by the vector $\bm{c}\coloneqq[c_{1},\ldots,c_{d}]^{T}$.
-With this re-parameterized problem, we can define the profile likelihood.
-If we fix $\bm{c}$ and $\bm{u}$, $\hat{\bm{\gamma}}$ can be analytically obtained in closed form:
-$\hat{\bm{\gamma}}=\left(D_{\bm{c}}B_{\bm{u}}^{T}B_{\bm{u}}D_{\bm{c}}+\lambda I\right)^{-1}D_{\bm{c}}B_{\bm{u}}^{T}\bm{y},$ (20)
-We can then get $\hat{\bm{\beta}}$ in closed form:
-$\hat{\bm{\beta}}=\bm{c}\circ\hat{\bm{\gamma}}\;,$ (21)
-where $\circ$” denoted element-wise multiplication.
-We can now substitute $\hat{\bm{\beta}}$ back into the original problem and take the constraint into consideration.
-This leads us to a Lagrangian minimization problem in the scaling variable $\bm{c}$ and neural network parameters $\bm{u}$:
-$\min_{\bm{c},\bm{u}}\|\bm{y}-B_{\bm{u}}\hat{\bm{\beta}}\|^{2}+\tilde{\mu}\sum_{i=1}^{d}\|\bm{c}\|^{2},$ (22)
-and we can optimize it through gradient descent. The training algorithm for this re-parameterized adaptive ridge regression is summarized in Alg. 1.
-Algorithm 1 Training of Re-parameterized Adaptive Ridge Regression
-0:  Training data $X\in\mathbb{R}^{n\times d}$, training target $\bm{y}\in\mathbb{R}^{n}$, neural network $g_{\bm{u}}(\cdot)$, regularization parameter $\lambda$, learning rate $\eta$, maximum iterations $T$.
-1:  Initialize network parameters $\bm{u}$ and scaling coefficients $\bm{c}$ (with $c_{i}>0$).
-2:  for $t=1,\dots,T$ do
-3:     $B_{\bm{u}}=X+X\circ G_{\bm{u}}$ // Compute modified design matrix
-4:     $D_{\bm{c}}=\operatorname{diag}(\bm{c})$ // Form diagonal scaling matrix
-5:     $\hat{\bm{\gamma}}=\bigl{(}D_{\bm{c}}B_{\bm{u}}^{T}B_{\bm{u}}D_{\bm{c}}+\lambda I\bigr{)}^{-1}D_{\bm{c}}B_{\bm{u}}^{T}\bm{y}$ // Solve for $\hat{\bm{\gamma}}$
-6:     $\hat{\bm{\beta}}=\bm{c}\circ\hat{\bm{\gamma}}$ // Compute regression coefficients
-7:     $L=\|\bm{y}-B_{\bm{u}}\hat{\bm{\beta}}\|^{2}+\tilde{\mu}\sum_{i=1}^{d}\|\bm{c}\|^{2}$ // Compute loss function
-8:     $\bm{c}\leftarrow\bm{c}-\eta\nabla_{\bm{c}}L,\quad\bm{u}\leftarrow\bm{u}-\eta\nabla_{\bm{u}}L$ // Gradient descent updates
-9:  end for
-10:  Return $\hat{\bm{\beta}},\bm{c},\bm{u}$ // Output final parameters
-B.2 Logistic regression: Iteratively reweighted least squares
-The procedure followed in the previous section can be extended to generalized linear models.
-Here we showcase how to do it for logistic regression.
-Logistic regression is a classification model that describes the probability that a binary outcome $y\in\{0,1\}$ occurs given a feature vector $\bm{x}\in\mathbb{R}^{d}$:
-$P(y=1|\bm{x})=\sigma(\bm{x}^{T}\bm{\beta})$ (23)
-where $\sigma(z)=\frac{1}{1+e^{-z}}$ is the sigmoid function.
-The model can then be trained by minimizing the negative log-likelihood.
-Similarly to the regression case discussed in the previous section, we add an adaptive ridge regression penalty to the objective:
-$\mathcal{L}(\bm{\beta})=-\sum_{i=1}^{n}[y_{i}\log\sigma(\bm{x}_{i}^{T}\bm{\beta})+(1-y_{i})\log(1-\sigma(\bm{x}_{i}^{T}\bm{\beta}))]+\sum_{j=1}^{d}\nu_{j}\beta_{j}^{2}\quad\text{ s.
-t.
-}\quad\sum_{j=1}^{d}\frac{1}{\nu_{j}}=\frac{d}{\lambda}$ (24)
-Now we can use the same re-parameterization used in the previous section for regression. We get the following optimization problem:
-$\begin{gathered}\min_{\bm{\gamma}}-\sum_{i=1}^{n}\left[y_{i}\log\sigma\left(\sum_{j=1}^{p}x_{ij}c_{j}\gamma_{j}\right)+(1-y_{i})\log\left(1-\sigma\left(\sum_{j=1}^{p}x_{ij}c_{j}\gamma_{j}\right)\right)\right]+\lambda\|\bm{\gamma}\|^{2}\\ \text{ s.t.
-}\quad\sum_{j=1}^{d}c_{j}^{2}=d,\quad c_{i}>0\end{gathered}$ (25)
-With the reparametrized objective we can consider the profile likelihood by fixing $\bm{c}$ and update $\bm{\gamma}$.
-To update $\bm{\gamma}$ we use the iteratively reweighted least squares (IRLS) algorithm.
-The associated update reads as:
-$\bm{\gamma}^{(t+1)}=(\tilde{X}^{T}W^{(t)}\tilde{X}+\lambda I)^{-1}\tilde{X}^{T}W^{(t)}\mathbf{z}^{(t)}$ (26)
-where $\tilde{X}$ is the modified design matrix with entries $\tilde{x}_{ij}=c_{j}x_{ij}$, $W^{(t)}$ is a diagonal weight matrix with entries $w_{i}^{(t)}=p_{i}^{(t)}(1-p_{i}^{(t)}),p_{i}^{(t)}=\frac{1}{1+e^{-\tilde{\bm{x}}_{i}^{T}\bm{\gamma}^{(t)}}}$, and $\mathbf{z}^{(t)}$ is the the vector of working response with entries $z_{i}^{(t)}=\bm{x}^{T}\bm{\beta}+\frac{y_{i}-p_{i}^{(t)}}{w_{i}^{(t)}}$.
-Note that in our model, our design matrix $X$ contains the nonlinear term as well so we would need to replace $X$ with $B_{\bm{u}}=X+X\circ G_{\bm{u}}$.
-We report the detailed optimization algorithm for the logistic regression case in Alg.
-2.
-Algorithm 2 IRLS Training for Logistic Regression
-0:  Training data $X\in\mathbb{R}^{n\times d}$, target labels $\bm{y}\in\{0,1\}^{n}$, neural network $g_{\bm{u}}(\cdot)$, regularization parameters $\lambda$, learning rate $\eta$, maximum iterations $T$.
-1:  Initialize parameters $\bm{u}$, scaling coefficients $\bm{c}$ (with $c_{i}>0$), and regression coefficients $\bm{\beta}$.
-2:  for $t=1,\dots,T$ do
-3:     $B_{\bm{u}}=X+X\circ G_{\bm{u}}$ // Compute transformed design matrix
-4:     $D_{\bm{c}}=\operatorname{diag}(\bm{c})$ // Form diagonal scaling matrix
-5:     $\tilde{X}=B_{\bm{u}}D_{\bm{c}}$
-6:     $\hat{\bm{y}}=B_{\bm{u}}\bm{\beta}$ // Compute logits
-7:     $\bm{\pi}=\sigma(\hat{\bm{y}})$
-8:     $W=\operatorname{diag}(\bm{\pi}\circ(1-\bm{\pi}))$
-9:     $\mathbf{z}=\hat{\bm{y}}+(\bm{y}-\bm{\pi})/\operatorname{diag}(W)$
-10:     $\bm{\gamma}=(\tilde{X}^{T}W\tilde{X}+\lambda I)^{-1}\tilde{X}^{T}W\mathbf{z}$
-11:     $\bm{\beta}=\bm{c}\circ\hat{\bm{\gamma}}$ // Update $\bm{\beta}$
-12:     $L=\operatorname{BCE}(\hat{\bm{y}},\bm{y})+\lambda\sum_{i}c_{i}^{2}$ // Compute total loss
-13:     $\bm{c}\leftarrow\bm{c}-\eta\nabla_{\bm{c}}L,\quad\bm{u}\leftarrow\bm{u}-\eta\nabla_{\bm{u}}L$ // Gradient descent step
-14:  end for
-15:  Return $\bm{\beta},\bm{c},\bm{u}$
-B.3 Extension to sub-$\ell_{1}$ pseudo-norms
-In some cases sparsity is a crucial aspect in the learning process.
-One of the limitations of lasso is the well-known over-shrinkage effect, which means that shrinkage of uninformative features could be achieved at the expense of shrinking the retained features as well.
-One way to avoid this is to use sub-$\ell_{1}$ pseudo-norms, which in the limist of the $\ell_{0}$ pseudo-norm achieve sparsity without any shrinkage.
-The main challenge with sub-$\ell_{1}$ pseudo-norm is that the resulting objective function becomes non-convex and harder to optimize.
-However, recent work has successfully used sub-$\ell_{1}$ norms as a prior on regression problems [Negri et al., 2023], which is similar to our setting.
-We now show how to use a sub-$\ell_{1}$ prior in our Adaptive Ridge regression objective, which we report below for completeness:
-$\min_{\bm{\beta}}\|X\bm{\beta}-\bm{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}\quad\text{s.t.}\quad\sum_{i=1}^{d}\frac{1}{\nu_{i}}=\frac{d}{\lambda}$ (27)
-In order to generalize to sub-$\ell_{1}$ pseudo-norms we need the following simple modification on $\nu_{i}$:
-$\sum_{i=1}^{d}\frac{1}{\nu^{\delta}}=C,\quad 0<\delta\leq 1.$ (28)
-We now showcase that we can follow a procedure completely similar to what we had with the original adaptive ridge regression.
-Fist, we reformulate the objective with the method of Lagrangian multipliers:
-$\mathcal{L}(\bm{\nu},\mu)=\|X\bm{\beta}-\bm{y}\|^{2}+\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}+\mu\sum_{i=1}^{d}\frac{1}{\nu_{i}^{\delta}}$ (29)
-Then, we find stationary points for $\nu_{i}$ by taking the gradients and setting them to zero:
-$\frac{\partial\mathcal{L}(\bm{\nu},\mu)}{\partial\nu_{i}}=\beta_{i}^{2}-\mu\delta\nu_{i}^{-(\delta+1)}=0\quad\Rightarrow$ $\quad\nu_{i}^{\delta+1}=\frac{\mu\delta}{\beta_{i}^{2}}$ (30)
-$\Rightarrow$ $\quad\nu_{i}=\left(\frac{\mu\delta}{\beta_{i}^{2}}\right)^{\frac{1}{\delta+1}}$
-The penalty in the original adaptive ridge regression can be rewritten as:
-$\sum_{i=1}^{d}\nu_{i}\beta_{i}^{2}=\sum_{i=1}^{d}\left(\frac{\mu\delta}{\beta_{i}^{2}}\right)^{\frac{1}{\delta+1}}\beta_{i}^{2}\propto\sum_{i=1}\beta_{i}^{2(1-\frac{1}{\delta+1})}.$ (31)
-Note that when $\delta=1$, we recover the Lasso $\ell_{1}$ norm.
-When $\delta\rightarrow 0$, it approaches the $\ell_{0}$ pseudo-norm, which basically counts the number of nonzero entries.
-We now apply the same re-parameterization:
-$\gamma_{i}=(\nu_{i}/\lambda)^{1/2}\beta_{i},\quad c_{i}=(\lambda/\nu_{i})^{1/2}\;.$ (32)
-The objective function then reads as
-$\min_{\bm{\gamma}}\|\bm{y}-XD_{\bm{c}}\bm{\gamma}\|^{2}+\lambda\|\bm{\gamma}\|^{2}\quad\text{s.t.}\quad\sum_{i=1}^{d}c_{i}^{2\delta}=\lambda^{\delta}C,\quad c_{i}>0\;.$ (33)
-For each fixed $\bm{c}$, $\bm{\gamma}$ can now be obtained in closed form:
-$\hat{\bm{\gamma}}=\left(D_{\bm{c}}X^{T}XD_{\bm{c}}+\lambda I\right)^{-1}D_{\bm{c}}X^{T}\bm{y},$ (34)
-and $\bm{\beta}$ can be computed as before by:
-$\hat{\bm{\beta}}=\bm{c}\circ\hat{\bm{\gamma}}$ (35)
-Once we compute the $\hat{\bm{\beta}}$, we can substitute it back into the original objective.
-Taking the constraint into consideration leads us to a Lagrangian minimization problem in the scaling variable $\bm{c}$ only:
-$\min_{\bm{c}}\|\bm{y}-X\hat{\bm{\beta}}\|^{2}+\tilde{\mu}\sum_{i=1}^{d}c_{i}^{2\delta},$ (36)
-and we can optimize it through gradient descent.
-Appendix C NIMO: Implementation details
-C.1 Model Formulation
-We formulate NIMO as follow:
-$f(\bm{x})=\beta_{0}+\sum_{j=1}^{d}x_{j}\beta_{j}(1+g_{\bm{u}_{j}}(\bm{x}_{-j}))\quad\text{s.t.}\quad g_{\bm{u}_{j}}(\mathbf{0})=0\;.$ (37)
-The architecture of our model involves two parts, as shown in Fig.
-2.
-The left part is the nonlinear neural network, which operates on the input feature $\bm{x}$ element-wise.
-The right part is the linear model, with coefficients $\beta_{j}$ multiplied by the neural network output and then summed together.
-In practice, each feature component $x_{j}$ of the input vector $\bm{x}$ is temporarily masked out to construct $\bm{x}_{-j}$, which is then passed through a neural network $g_{\bm{u}_{j}}$.
-The output of this network is added to a constant $1$ and then multiplied by the original feature value $x_{j}$, effectively producing a rescaled version of that feature:
-$\tilde{x}_{j}=x_{j}\cdot(1+g_{\bm{u}_{j}}(\bm{x}_{-j}))$ (38)
-This rescaling can be interpreted as an element-wise modulation of the input features based on context.
-The rescaled features are then linearly combined using fixed coefficients $\bm{\beta}$ to produce the final prediction:
-$y=\sum_{j=1}^{d}\tilde{x}_{j}\beta_{j}.$
-C.2 Practical Implementation
-We implement our model with PyTorch Lightning [Falcon and The PyTorch Lightning team, 2019].
-While in principle each feature $x_{j}$ could have its own dedicated network $g_{\bm{u}_{j}}$, this can be computationally expensive for high-dimensional inputs.
-To reduce overhead, we instead use a shared neural network $g_{\bm{u}}$ across all components.
-Positional Encoding
-To ensure the network remains aware of which feature is currently masked out, we append a positional encoding to $\bm{x}_{-j}$ before passing it through the shared network.
-We assign each masked input feature $\bm{x}_{-j}$ a unique binary vector corresponding to its index $j$, represented in fixed-length binary form using $\lfloor\log_{2}d\rfloor+1$ bits.
-This compact encoding allows the shared neural network to remain index-aware with minimal overhead.
-Group Penalty and Noise Injection
-The shared network is implemented with 3 fully-connected (fc) layers.
-To encourage the network to select only certain relevant features contributing to the nonlinearity, we apply a group penalty on the first fc layer.
-However, due to the compensation effect from the subsequent layers, the sparsity is difficult to achieve by simply applying this penalty.
-To mitigate this issue, we inject noise into the output of the first layer by adding scaled Gaussian noise.
-Concretely, we draw noise from a normal distribution with mean 0 and variance 1, scale it by 0.2, and add it to the output.
-The noise-injected output is then passed through a tanh activation function before being fed into the subsequent layers.
+As a result, the network is encouraged to focus on strong and robust signals associated with informative input features.
+This approach works very well in practice and allows to easily achieve sparsity in the neural network.
 Other Details
-After the second fc layer, a sin activation function is applied to capture potentially repeating or high-frequency cyclic patterns in the data [Sitzmann et al., 2020].
+After the second fc layer, a sin activation function is applied to capture potentially repeating or high-frequency cyclic patterns in the data (Sitzmann et al., 2020).
 The learning rate is 1e-3 for synthetic experiments, and 5e-3 for real-world datasets.
-We choose Adam [Kingma, 2014] as the default optimizer.
-For regression, we regularize the network also by restricting the output range within $[-1,1]$, which works well in practice.
-For classification, we use a less restrictive reduction to the range $[-1,3]$.
+We choose Adam (Kingma, 2014) as the default optimizer.
+In all experiments, the lasso and group penalties are selected by performing a grid search over a predefined set of candidate values.
 Appendix D Experiments
 We conduct experiments on both synthetic and real-world datasets.
-All experiments are conducted on a workstation equipped with a single user-grade NVIDIA GeForce RTX 2080 Ti GPU.
-The methods we compare with are linear model (Lasso for regression, Logistic Regression for classification), neural network, LassoNet [Lemhadri et al., 2021] and Contextual Lasso [Thompson et al., 2023] (for regression only).
-Here we provide the details about the experiment setups.
-D.1 Implementation details / Hyperparameters
+All experiments are conducted on a workstation equipped with a single user-grade NVIDIA GeForce RTX 5080 GPU.
+The methods we compare with are linear model (Lasso for regression, Logistic Regression for classification), a vanilla neural network (NN(Goodfellow et al., 2016), and several state of the art interpretable approaches, namely LassoNet (Lemhadri et al., 2021), Neural Additive Models (NAMs(Agarwal et al., 2021) and Interpretable Mesomorphic Networks (IMNs(Kadra et al., 2024).
+D.1 Methods implementation
 Linear Model
-We use scikit-learn [Pedregosa et al., 2011] to implement the Lasso and Logistic regression model.
+We use scikit-learn (Pedregosa et al., 2011) to implement the Lasso and Logistic regression model.
 More specifically, we use LassoCV and LogisticRegressionCV with default 5-fold cross validation to select the best model.
@@
 Neural Network
-We implement the naive neural network (MLP) with PyTorch Lightning [Falcon and The PyTorch Lightning team, 2019].
-For a fair comparison, the MLP has the same structure as the nonlinear part in our NIMO model, except no positional encoding.
-We use Adam [Kingma, 2014] as the default optimizer.
+We implement the naive neural network (NN) with PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019).
+For a fair comparison, the NN has the same structure as the nonlinear part in our NIMO model, except no positional encoding.
+We use Adam (Kingma, 2014) as the default optimizer.
 The learning rate is 1e-3 for synthetic experiments, and 5e-3 for real-world datasets.
@@
 The one with the smallest validation loss is chosen.
-Contextual Lasso
-We get the code of Contextual Lasso from the original Github repository  [^3] https://github.com/ryan-thompson/ContextualLasso.jl, which was implemented in Julia language.
-We follow the examples in the documentation and configure the model to have a structure and number of parameters comparable to our own model.
+NAM
+The official NAM implementation was in Tensorflow.
+We chose a well-maintained PyTorch implementation from the GitHub repository  [^3] https://github.com/lemeln/nam.
+This implementation provides a scikit-learn style interface, which can be easily used.
+IMN
+We use the original PyTorch implementation of IMN from their GitHub repository  [^4] https://github.com/ArlindKadra/IMN.
+All the arguments are kept at their default values except modifying the data loading logic to fit our datasets.
+For all the methods above, we tried our best to make sure that they achieve reasonably good and consistent performance on all the datasets.
 D.2 Synthetic experiments: Setup
@@
 Random Gaussian noises are added to the targets.
-The feature values are sampled uniformly between $-2$ and $2$.
-For logistic regression, we pass the targets through a logistic (sigmoid) function to get probabilities between 00 and $1$, and then sample the final binary labels using a Bernoulli distribution with those probabilities.
+The feature values are sampled uniformly between 2-2 and $2$.
+For logistic regression, we pass the targets through a logistic (sigmoid) function to get probabilities between 0 and $1$, and then sample the final binary labels using a Bernoulli distribution with those probabilities.
 In all the settings, we use $200$ training samples and $100$ samples each for validation and test.
@@
   • Setting 0 (toy example), $n=200,p=3$
-    $y=$ $+3\cdot x_{1}\cdot[1+\tanh(10x_{2})]$ (39)
+[…]
     $-3\cdot x_{2}\cdot\left[1+\sin(-2x_{1})\right]$
@@
     $+\epsilon$
-  • Setting 4, $n=200,p=10$
+  • Setting 4 (vanilla regression), $n=200,p=10$
     $y$ $=-5\cdot x_{1}-4\cdot x_{2}-3\cdot x_{3}-\dots+0\cdot x_{6}$ (43)
@@
 Settings for classification
+In order to create data for synthetic classification we follow a similar procedure as in the regression cases, where we directly control the linear coefficients and the nonlinear interactions.
+The only difference is that we further apply a link function and binarize the output and get the labels.
+Note that this way we loose information about the original linear coefficients, which cannot be recovered exactly anymore.
+The sparsity in the coefficients still remains crucial.
 For all the settings, we generate $y$ first, and then generate labels through the following procedure:
@@
   • Setting 2, $n=200,p=10$
-    $y=$ $+10\cdot x_{1}\cdot[1+2\cdot\tanh(2x_{2})+\sin(x_{4})]$ (45)
-    $+20\cdot x_{2}\cdot[1+2\cdot\cos(2x_{1})]$
-    $-20\cdot x_{3}\cdot[1+2\cdot\arctan(x_{2}x_{4})]$
-    $+10\cdot x_{4}$
-    $-10+\epsilon$
+[…]
+[…]
+[…]
+[…]
+[…]
   • Setting 3, $n=200,p=50$
-    $y=$ $-20\cdot x_{1}\cdot[1+\tanh(x_{2}x_{4})]$ (46)
-    $+20\cdot x_{2}\cdot\left[1+\frac{2}{\pi}\arctan(x_{4}-x_{5})\right]$
-    $+30\cdot x_{4}\cdot\left[1+\tanh(x_{2}+\sin(x_{5}))\right]$
-    $-10\cdot x_{5}\cdot\left[1+(2\sigma(x_{1}x_{4})-1)\right]$
+[…]
+[…]
+[…]
+[…]
     $+\epsilon$
@@
 Below, we present additional experimental results for the synthetic settings described above.
+Regression
+As we mentioned in the main paper, we further apply group $\ell_{2}$ regularization to the weight matrix of the first fully connected (fc) layer, encouraging feature-level sparsity and clarifying how inputs contribute to predictions.
+We provide the details in Appendix C.2.
+Here, we show the sparse features selected by the neural network in Figure 8.
+We observe that even though all features are input to the neural network, only a small subset of features are selected.
+Moreover, these selected features align with those involved in the nonlinearity computation in our settings, which demonstrates the model’s ability to effectively identify and focus on the most relevant features for the task.
 Figure 8: Sparsity and feature selection in the first fc layer of NIMO for regression settings. Left: Setting 1. Middle: Setting 2. Right: Setting 3.
-Regression
-The advantage of our model is that it not only recovers the underlying linear coefficients but also selects sparse features for the neural network, thanks to the noise injection and group penalty.
-In the main paper, we presented the recovered linear coefficients.
-Here, we show the sparse features selected by the neural network in Figure 7.
-We observe that even though all features are input to the neural network, only a small subset of features is selected.
-Moreover, these selected features align with those involved in the nonlinearity computation in our settings, which demonstrates the model’s ability to effectively identify and focus on the most relevant features for the task.
 Classification
+Our method can be extended to generalized linear models and here we showcase it for logistic regression.
+Also in this case we show various settings for different dimensionality, nonlinear terms and different number of uninformative features.
+In Table 3 we report the classification accuracy on the test set for all methods.
+Overall, our model significantly outperforms logistic regression on all settings.
+When analyzing the learned coefficient, we can see that only NIMO recovers the correct sparsity patterns.
+As argued before, this allows to learn the correct nonlinear interactions.
+Other nonlinear models and hybrid approaches perform similarly to NIMO but fail to provide the interpretability of the coefficients as in a linear model.
+Table 3: Classification accuracy for synthetic classification settings. We compare with logistic regression, NN and LassoNet. In each setting we use 200 samples.
+Features Log. Regr. NN NIMO LassoNet
+Setting 1 3 0.59 0.90 0.92 0.89
+Setting 2 10 0.68 0.83 0.85 0.82
+Setting 3 50 0.74 0.69 0.84 0.90
 In classification, the presence of the link function can obscure the underlying signal, so while a linear decision boundary may suffice, recovering the true generative coefficients is more difficult than in regression, where the model directly learns to approximate the target surface.
@@
 D.4 Real-dataset experiments
-We conduct experiments on two real-world datasets: the diabetes dataset [Efron et al., 2004] and the Boston housing dataset [Belsley et al., 2005].
+We conduct experiments on three real-world datasets: the diabetes dataset (Efron et al., 2004), the Boston housing dataset (Belsley et al., 2005) and the superconductivity dataset (Hamidieh, 2018).
 Each dataset is randomly split into training (60%), validation (20%), and test (20%) sets, repeated over 5 independent runs.
@@
 Error bars and confidence intervals are reported across these different splits.
+The main results are shown in Figure 5.
+Diabetes dataset
+The diabetes dataset is a classic regression dataset consisting of 442 patients, each with 10 numeric features, and the target measures disease progression one year after baseline.
+As shown in Figure 11 (left), we observe that all methods achieve similar MSE losses and the neural network (also the hybird methods) shows no clear advantage on this dataset.
+Since the diabetes dataset is commonly used for linear regression tasks, it is not surprising that the data does not exhibit complex nonlinear interactions among features.
+We can verify this claim by looking at the weight sparsity of the first fc layer in NIMO, shown in Figure 11 (right).
+Relevantly, NIMO achieves a significantly sparser solution than Lasso regression, see Figure 11 (middle).
+This is particularly evident by looking at the coefficients associated with the feature “s1”, “s2” and “s4”.
+Figure 11: Diabetes datasets results.
+Left: MSE loss on test set.
+All methods achieve comparable results.
+Middle: Learned linear coefficients by Lasso and NIMO.
+NIMO selects sparser coefficients (“s1, “s2”, “s4”).
+Right: the first fc layer of NIMO is very sparse; few nonlinearities are relevant.
+Boston dataset
+The Boston housing dataset contains 506 instances, each with 13 features, and the target is the median value of owner-occupied homes.
+As shown in Figure 12 (left), our method performs significantly better than Lasso regression and shows compatible MSE test loss with the other nonlinear and hybrid approaches.
+This suggests that nonlinear interactions are crucial to explain the data.
+While the other methods are able to capture those nonlinearities, they lack the interpretability provided by our model.
+It is important to note that the Boston housing dataset has been criticized for containing controversial features  [^5] https://fairlearn.org/main/user_guide/datasets/boston_housing_data.html.
+The feature of interest is termed “B”.
+From the coefficient plot in Figure 12 (middle), we observe that the significance of the feature “B” is greatly reduced in our model compared to Lasso regression.
+This surprising result suggests that the role of the feature “B” towards the prediction might be irrelevant.
+Note that the feature “B” is also not relevant for the nonlinear terms, as shown in Figure 12 (right).
+Figure 12: Boston datasets results.
+Left: Comparison of MSE losses.
+NIMO performs on par with other methods and significantly outperforms Lasso, suggesting nonlinear interactions are crucial.
+Middle: Learned linear coefficients by Lasso and NIMO.
+The two models learn coefficient with a comparable sparsity.
+Relevantly, the feature “B” is almost irrelevant for NIMO.
+Right: sparsity in the first fc layer of NIMO.
+Superconductivity
+The superconductivity dataset contains 21,263 instances, each with 81 features. The target is to predict the critical temperature based on the features extracted.
+Figure 13: Superconductivity dataset results. Left: Comparison of MSE losses. Right: sparsity in the first fc layer of NIMO.
+In Figure 13, we observe that NIMO achieves the best performance.
+With respect to sparsity in the first fc layer, the presence of strong correlations among several of the 81 features makes it challenging to obtain clear sparsity patterns.
+Nevertheless, certain patterns can still be discerned (Figure 13, right).
+Figure 14 shows the coefficients for Lasso and NIMO on this dataset.
+Figure 14: Coefficients for Lasso and NIMO on superconductivity dataset.
+D.5 Ablation study of feature masking
+To verify the effectiveness of masking out the $j$-th component of an instance when passing it through the neural network, we conducted an ablation study.
+Specifically, we examined the effect of masking out the $j$-th component as well as the effect of removing position encoding.
+From the ablation results across the three settings (Figure 7, Figure 15, and Figure 16), we observe several consistent patterns:
+  • Without masking the $j$-th component, the first fc layer of the neural network can still exhibit sparsity patterns, but they are less distinct than when masking is applied.
+We attribute this to the position encoding, which enables the network to identify the feature index.
+  • Without masking the $j$-th component, the magnitudes of the linear coefficients are distorted, reducing the interpretability of the marginal effects at the mean (MEM).
+  • Further removing the position encoding distorts the sparsity patterns in the first fc layer and substantially alters the magnitude of the linear coefficients, thereby undermining interpretability.
+Figure 15: The effect of feature masking and position encoding on the linear coefficients and the sparsity of the first fc layer on setting 2.
+Figure 16: The effect of feature masking and position encoding on the linear coefficients and the sparsity of the first fc layer on setting 3. Coefficients are displayed for the first 10 features.
+Appendix E LLM statement
+ChatGPT is used only for polishing the sentences.

### v3

@@
 However, neural networks in general lack interpretability of the predictions in terms of the input features, hence the term “black-box” models.
-In contrast, classical models such as linear regression and decision trees are often considered as highly interpretable because of their transparent structure and simple decision rules, but they can lack predictive power on complex, high-dimensional, or highly nonlinear tasks (Hastie, 2009; Breiman et al., 2017).
+In contrast, classical models such as linear regression and decision trees are often considered as highly interpretable because of their transparent structure and simple decision rules, but they can lack predictive power on complex, high-dimensional, or highly nonlinear tasks (Hastie et al., 2009; Breiman et al., 2017).
 This tension between accuracy and interpretability has motivated much recent work on models and methods that combine the expressive power of neural networks with the transparency of simpler models (Lemhadri et al., 2021; Thompson et al., 2023; Wu et al., 2017; 2021).
-Meanwhile, many model-agnostic, post-hoc explainers have been proposed, such as SHAP (Lundberg & Lee, 2017) and LIME (Ribeiro et al., 2016), providing instance-level feature attributions for arbitrary predictors.
+Meanwhile, many model-agnostic, post-hoc explainers have been proposed, such as SHAP (Lundberg and Lee, 2017) and LIME (Ribeiro et al., 2016), providing instance-level feature attributions for arbitrary predictors.
 But these post-hoc explanations are approximations and may depend on the choices of parameters, such as the background distribution and kernel width.
@@
 To further enhance interpretability, we impose sparsity by formulating the problem as adaptive ridge regression (Grandvalet, 1998).
-In the network, we further apply group $\ell_{2}$ regularization (Yuan & Lin, 2006) to the weight matrix of the first fully connected layer, encouraging feature-level sparsity and clarifying how inputs contribute to predictions.
+In the network, we further apply group $\ell_{2}$ regularization (Yuan and Lin, 2006) to the weight matrix of the first fully connected layer, encouraging feature-level sparsity and clarifying how inputs contribute to predictions.
 Overall, the contributions of the present work can be summarized as follows:
@@
 Most of the methods, especially in deep learning, provide post-hoc explanations.
-The most popular ones include LIME (Ribeiro et al., 2016), SHAP (Lundberg & Lee, 2017) and Grad-CAM (Selvaraju et al., 2020).
+The most popular ones include LIME (Ribeiro et al., 2016), SHAP (Lundberg and Lee, 2017) and Grad-CAM (Selvaraju et al., 2020).
 On the other hand, interpretability focuses on understanding the inner mechanisms and decision processes of the model, and most of the methods that claimed to be interpretable involve inherently understandable models, such as decision trees (Wu et al., 2017; 2021), Lasso (Tibshirani, 1996) and Explainable Boosting Machines (Lou et al., 2013).
@@
 Several works have proposed combining linear models with neural networks to improve interpretability.
-Neural Additive Models (NAMs) (Agarwal et al., 2021) extend generalized additive models by learning one neural network per feature, whose outputs are summed to form the prediction.
+Neural Additive Models (NAMs) (Agarwal et al., 2021) extend generalized additive models (GAMs) by learning one neural network per feature, whose outputs are summed to form the prediction.
 This design preserves interpretability at the feature level but does not have the ability to capture feature interactions.
+NODE-GAM (Chang et al., 2021) is also a neural generalized additive model that uses differentiable tree ensembles to learn interpretable univariate and pairwise shape functions.
+However, similar to NAM, both of their interpretability comes from the shape function, which is input dependent, and cannot provide a global population-level summary.
 LassoNet (Lemhadri et al., 2021) integrates a neural network with Lasso regression by enforcing that a feature can be used in the nonlinear part only if its corresponding linear coefficient is nonzero.
@@
 While this increases flexibility, it sacrifices global interpretability of the coefficients, making it difficult to recover clear baseline effects.
+Interpretable Mixture of Experts (IME) (Ismail et al., 2022) routes each sample to an interpretable expert (e.g., a linear model), ensuring that for that sample the entire decision process is an exact and transparent explanation of which expert was chosen and why.
+While IME can maintain high accuracy and provide faithful local explanations, its global interpretability remains limited.
 In contrast, our proposed model (NIMO) preserves the global interpretability of linear coefficients while augmenting them with nonlinear corrections.
@@
 In contrast, for NIMO, the MEMs directly correspond to the linear coefficients $\bm{\beta}$, capturing the additive population-level contribution of each feature.
+While MEM is widely used as a global effect measure in statistics, it is important to acknowledge its general limitations, as discussed in Scholbeck et al.
+(2024).
+MEM is meaningful and interpretable in models that include an explicit global linear component, but for highly nonlinear functions the MEM may fail to reflect the true relevance of individual features.
+However, NIMO is explicitly designed to avoid such cases: its architecture explicitly separates a global linear component (captured by $\beta$) from local, per-instance nonlinear adjustments.
+This structural constraint ensures that the global linear contribution remains identifiable and stable.
+Consequently, MEM remains appropriate and interpretable within NIMO’s model class, even though it may be unsuitable as a universal feature-importance metric for fully unconstrained nonlinear functions.
 3.3 Training via parameter elimination
@@
 Let us first consider the standard ridge regression setting: $\mathcal{L}(\bm{\beta},\bm{u})=\lVert\mathbf{y}-B_{\bm{u}}\bm{\beta}\rVert^{2}+\lambda\lVert\bm{\beta}\rVert^{2}$.
-To separate the optimization over $\bm{\beta}$ and $\bm{u}$, we take inspiration from the profile likelihood approach (Venzon & Moolgavkar, 1988; Murphy & Van der Vaart, 2000).
+To separate the optimization over $\bm{\beta}$ and $\bm{u}$, we take inspiration from the profile likelihood approach (Venzon and Moolgavkar, 1988; Murphy and Van der Vaart, 2000).
 The key idea is to eliminate $\bm{\beta}$ by first solving for its closed-form expression in terms of $\bm{u}$, and then substituting it back into the objective.
@@
 The full procedure is summarized in Algorithm 1 in the Appendix.
-A well-known limitation of Lasso is the over-shrinkage effect, where large coefficients are excessively penalized (Fan & Li, 2001).
+A well-known limitation of Lasso is the over-shrinkage effect, where large coefficients are excessively penalized (Fan and Li, 2001).
 In some applications, this can lead to significant bias and degrade model performance.
@@
 As before, we first get a closed-form expression for $\bm{\beta}$ in terms of $\bm{u}$ and then substitute it back to optimize over the neural network parameters $\bm{u}$.
-What is different is that in each optimization step we use the iteratively reweighted least squares (IRLS) surrogate (Green, 1984; McCullagh & Nelder, 1989), which is a weighted least-squares approximation of the original problem:
+What is different is that in each optimization step we use the iteratively reweighted least squares (IRLS) surrogate (Green, 1984; McCullagh and Nelder, 1989), which is a weighted least-squares approximation of the original problem:
 $\min_{\bm{\beta},\bm{u}}\;\tfrac{1}{2}\bigl\|W^{1/2}(\mathbf{z}-B_{\bm{u}}\bm{\beta})\bigr\|^{2}+\lambda\|\bm{\beta}\|_{1}\;,$ (7)
@@
 In this paper, we compare our method against Lasso (Tibshirani, 1996), a vanilla neural network (NN(Goodfellow et al., 2016), and several state of the art interpretable approaches, namely LassoNet (Lemhadri et al., 2021), Neural Additive Models (NAMs(Agarwal et al., 2021) and Interpretable Mesomorphic Networks (IMNs(Kadra et al., 2024).
-LassoNet encourages the network to use only a subset of the available input features by enforcing sparsity with a novel objective.
-NAMs learn a linear combination of neural networks that each attend to a single input feature.
-IMNs optimize deep hypernetworks to generate explainable linear models on a per-instance basis.
 Table 1: MSE loss for synthetic regression settings.
@@
 NIMO 0.024 0.197 0.380
+LassoNet encourages the network to use only a subset of the available input features by enforcing sparsity with a novel objective.
+NAMs learn a linear combination of neural networks that each attend to a single input feature.
+IMNs optimize deep hypernetworks to generate explainable linear models on a per-instance basis.
 Protocol:
@@
 Recall that MEM estimates marginal effects at the average values of the input features, reflecting the additive population-level contribution of each feature.
-The visualizations reveal that the MEM from NIMO differ substantially from those of Lasso, reflecting NIMO’s ability to capture nonlinear feature interactions that Lasso cannot represent.
-For the superconductivity dataset, which involves a much larger set of features, the corresponding MEM visualization is provided in the Appendix.
+We observe that the MEM profiles of Lasso and NIMO differ across several features.
+These differences do not necessarily imply the presence of nonlinear feature interactions; rather, both datasets contain groups of strongly correlated predictors, making feature selection inherently unstable.
+Consequently, Lasso and NIMO may highlight different features within the same correlated group, even if their predictive roles are similar.
+In Appendix D.8, we experimentally showcase the degree of nonlinearities in each dataset.
+For the superconductivity dataset, which involves a much larger set of features, the corresponding MEM visualization is provided in the Appendix in Figure 14.
 We then use MEM to rank the features on the diabetes and Boston housing datasets.
-For comparison, we also apply the post-hoc explainer SHAP (Lundberg & Lee, 2017) to our model and rank features according to the global SHAP values.
+For comparison, we also apply the post-hoc explainer SHAP (Lundberg and Lee, 2017) to our model and rank features according to the global SHAP values.
 The results are shown in Table 2.
@@
 Figure 7: The effect of feature masking and position encoding on the linear coefficients and the sparsity of the first fc layer on setting 1.
+4.4 Limitations
+The proposed approach is designed to preserve the global, population-level interpretability inherited from linear models while simultaneously learning nonlinear corrections to linear predictions through neural networks, thereby enhancing the expressive capacity of linear models.
+Thus, the underlying assumption is that the data can be well described by adding nonlinear refinements on top of a linear model.
+In settings dominated by strong self-interactions or highly nonlinear regimes, NIMO cannot fully capture the underlying relationships in an interpretable manner.
+Assume the data-generating process contains a self-interacting term of the form $x_{j}f(x_{j})$, where $f$ is a linear or nonlinear function (e.g.
+$x\sin(x)$ or $x^{2}$).
+To maintain interpretability of the linear coefficients, we imposed the constraint that $g_{\bm{u}_{j}}(\mathbf{\mathbf{x}_{-j}})$ does not depend on $x_{j}$, which prevents NIMO from learning such self-interacting nonlinearities.
+However, if such a feature is known to play an important role in the dataset, simple feature engineering can resolve the issue.
+To experimentally and explicitly illustrate both the limitation of NIMO and a straightforward remedy, we conduct an ablation study on a synthetic dataset.
+The data are generated as
+$y=x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+\epsilon.$ (9)
+Empirically, we demonstrate that naively applying NIMO to this dataset results in poor performance.
+However, after applying a polynomial basis expansion of degree 2, NIMO can perfectly solve the problem while still providing global, population-level interpretability through the linear coefficients $\bm{\beta}$.
+More detailed experimental results can be found in Appendix D.6.
 5 Conclusions
@@
 We demonstrate the effectiveness of the proposed approach on both synthetic and real datasets, showing competitive predictive performance while delivering meaningful and globally consistent interpretations.
-A key limitation of the approach is that it assumes the data can be described as a linear model with nonlinearity corrections; highly nonlinear datasets or features with self-interactions (e.g., $x_{j}f(x_{j})$) may not be fully captured.
-In such cases, careful feature engineering may be required to preserve interpretability.
 Appendix
 Appendix A Equivalence of Adaptive Ridge regression and Lasso
-The equivalence in the solution of Adaptive ridge regression and Lasso regression is a well-known fact (Tibshirani, 1996; Grandvalet, 1998; Grandvalet & Canu, 1998).
+The equivalence in the solution of Adaptive ridge regression and Lasso regression is a well-known fact (Tibshirani, 1996; Grandvalet, 1998; Grandvalet and Canu, 1998).
 Since both our model relies on this fact both in the regression and classification settings, we report its proof below.
@@
 where $\nu_{i}>0$ are the penalty parameters for each coefficient $\beta_{i}$, and $\lambda>0$ is a predefined parameter.
-To avoid divergent solutions we use the re-parameterization employed in Grandvalet (1998); Grandvalet & Canu (1998):
+To avoid divergent solutions we use the re-parameterization employed in Grandvalet (1998); Grandvalet and Canu (1998):
 $\gamma_{i}=(\nu_{i}/\lambda)^{1/2}\beta_{i},\quad c_{i}=(\lambda/\nu_{i})^{1/2}\;.$ (20)
@@
 C.2 Practical Implementation
-We implement our model with PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019).
+We implement our model with PyTorch Lightning (Falcon and The PyTorch Lightning team, 2019).
 While the model definition is simple, in practice, several strategies are needed to make the training more efficient and effective.
@@
 Neural Network
-We implement the naive neural network (NN) with PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019).
+We implement the naive neural network (NN) with PyTorch Lightning (Falcon and The PyTorch Lightning team, 2019).
 For a fair comparison, the NN has the same structure as the nonlinear part in our NIMO model, except no positional encoding.
@@
 The learning rate is 1e-3 for synthetic experiments, and 5e-3 for real-world datasets.
-Strong dropout regularization (p=0.6) is used to avoid overfitting.
+Strong dropout regularization (p=0.6) is used to avoid overfitting for small scale datasets (including synthetic datasets, diabetes and Bostong housing datasets).
+For the relatively large superconductivity dataset, small dropout rate (p=0.1) is used.
 LassoNet
@@
 Figure 13: Superconductivity dataset results. Left: Comparison of MSE losses. Right: sparsity in the first fc layer of NIMO.
-In Figure 13, we observe that NIMO achieves the best performance.
+In Figure 13, we observe that NIMO achieves the best performance comparable to general neural network.
 With respect to sparsity in the first fc layer, the presence of strong correlations among several of the 81 features makes it challenging to obtain clear sparsity patterns.
@@
 Figure 14 shows the coefficients for Lasso and NIMO on this dataset.
-Figure 14: Coefficients for Lasso and NIMO on superconductivity dataset.
+Figure 14: Coefficients (also MEMs) for Lasso and NIMO on superconductivity dataset.
 D.5 Ablation study of feature masking
@@
 Figure 16: The effect of feature masking and position encoding on the linear coefficients and the sparsity of the first fc layer on setting 3. Coefficients are displayed for the first 10 features.
+D.6 Limitations of NIMO
+To clearly and explicitly illustrate the limitations of our model, we conduct an additional set of ablation studies in this subsection.
+As discussed above, due to its architectural constraints, NIMO is not able to capture datasets with strong nonlinearity or purely self-interacting features.
+To demonstrate this, we create the following synthetic dataset:
+$y=x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+\epsilon,$ (48)
+where the feature values are sampled uniformly from [2,2][-2,2] and $\epsilon\sim\mathcal{N}(0,1)$ is a Gaussian noise.
+We compare the predictive performance of Lasso, NIMO, and LassoNet in Table 4.
+As expected, NIMO performs similarly poor to Lasso, while LassoNet is still able to model this dataset effectively.
+From Figure 17 (left), we observe that the coefficients learned by NIMO closely resemble those of Lasso, and NIMO effectively learns only a constant intercept $\beta_{0}$.
+In addition, we plot the weight matrix of the first fc layer of the neural network and observe that all weights are close to zero.
+This further confirms that, in this setting, NIMO reduces to a purely linear model without contributing additional nonlinear modeling capacity.
+This outcome is fully consistent with our expectations and illustrates a known limitation of NIMO: it cannot recover self-nonlinearities such as $x_{j}^{2}$ due to its interpretability-preserving architectural design.
+Although LassoNet is able to model this type of nonlinearity, it does not offer the same ability as NIMO or linear models to provide global, population-level interpretability through stable and directly interpretable learned coefficients.
+Table 4: Predictive performance on the synthetic self-interaction dataset.
+NIMO performs similarly to Lasso in the original feature space but succeeds after basis expansion, where quadratic terms are explicitly included.
+Method Original Basis Expansion
+(#features) (3) (9)
+Lasso 5.307 1.009
+LassoNet 1.027 1.012
+NIMO 5.340 1.009
+Figure 17: Learned coefficients and first-layer weights on the self-interaction dataset.
+(Left) In the original feature space, NIMO collapses to a constant model, matching Lasso.
+(Right) With basis expansion, NIMO recovers the quadratic effects.
+However, if we know or suspect that certain feature interactions play a role in the dataset, we can model them explicitly through feature engineering.
+In this experiment, we apply a second-order polynomial basis expansion, transforming the original feature space $x_{1},x_{2},x_{3}]$ into $x_{1},\,x_{2},\,x_{3},\,x_{1}^{2},\,x_{1}x_{2},\,x_{1}x_{3},\,x_{2}^{2},\,x_{2}x_{3},\,x_{3}^{2}]$.
+From Table 4, we can see that after basis expansion, both NIMO and Lasso achieve performance comparable to LassoNet.
+In Figure 17 (right), NIMO also correctly recovers the linear coefficients associated with the quadratic terms $x_{1}^{2}$, $x_{2}^{2}$, and $x_{3}^{2}$, demonstrating that NIMO can model self-nonlinearities once they are explicitly included in the feature representation.
+In summary, while NIMO cannot capture self-interactions in the raw feature space due to its interpretability-preserving architecture, simple basis expansions allow it to model such nonlinearities effectively while retaining global interpretability.
+D.7 Local Explanations
+NIMO provides intelligible global interpretability through the marginal effects at the mean (MEM) and per-instance adjustments via the correction networks.
+In the sections above, we have demonstrated the effectiveness of NIMO’s global interpretation.
+Here, we briefly showcase its ability to provide meaningful local, per-instance explanations.
+Local Explanation on Synthetic Dataset.
+For local explanations, NIMO defines a local coefficient for each feature,
+$s_{j}(\mathbf{x})=\beta_{j}\big(1+g_{\mathbf{u}_{j}}(\mathbf{x}_{-j})\big),$ (49)
+which represents how the global effect $\beta_{j}$ is modulated by the specific input context. Based on this, we compute the local contribution
+$c_{j}(\mathbf{x})=x_{j}\,s_{j}(\mathbf{x})$ (50)
+for each feature at each instance.
+The scatter plots in Figure 18 compare the ground-truth contribution term for feature $j$ with NIMO’s estimated local contribution $c_{j}(\mathbf{x})$ for synthetic setting 1.
+In this setting, the dataset contains five features, but only the first three have nonzero linear coefficients in the data-generating process.
+As shown in the figure, the first three features exhibit scatter points that align closely with the identity line, while the remaining two features produce scatter distributions centered near the origin.
+These results indicate that NIMO’s local explanations faithfully capture the underlying structure of the data.
+Figure 18: Local explanations for each feature on the synthetic setting 1.
+Local Explanation on Real Dataset.
+For real datasets, where ground-truth local contributions are unavailable, we compare NIMO’s local contributions $c_{j}(\mathbf{x})$ with SHAP values.
+As shown in Figure 19, the scatter plots for most features align closely with the identity line, indicating that NIMO’s local explanations are consistent with widely used post-hoc explanation tools while remaining intrinsic to the model.
+Due to the sparsity constraint, some coefficients learned by NIMO are exactly zero, which forces the corresponding local contributions to be zero as well.
+In such cases, the scatter plot reflects this behavior clearly.
+For example, the local contributions for the feature NOX collapse to zero for all instances.
+Figure 19: Local explanations for each feature on the Boston housing dataset.
+D.8 Comparison of the Degree of Nonlinearity
+In the main paper (Sec.
+4, Figure 5), we observed a substantial difference in predictive performance between the Diabetes and Boston Housing datasets, and attributed this phenomena to the relatively low level of nonlinearity present in the Diabetes dataset.
+To experimentally and quantitatively verify this claim, we compute the absolute magnitude of the nonlinear correction for each feature,
+$n_{j}(\mathbf{x})=\big|1+g_{\mathbf{u}_{j}}(\mathbf{x}_{-j})\big|,$ (51)
+and then average these values over all test samples for each feature.
+Figure 20 shows the resulting mean nonlinearity levels for both datasets.
+The Boston Housing dataset exhibits noticeably larger deviations from 1 across many features, indicating stronger nonlinear interactions.
+In contrast, the Diabetes dataset shows values much closer to 1 for all features, confirming that its underlying structure is primarily linear.
+These results support our interpretation of the performance differences observed in the main experiments.
+Figure 20: Comparison of the degree of nonlinearity in the Diabetes and Boston Housing datasets.
+D.9 Ablation study of the sparse penalty on the first FC layer
+To verify whether applying the group penalty to the first fc layer of the neural network is beneficial, we conduct an ablation study on the synthetic dataset (setting 1).
+Figure 21: Ablation study of the group penalty on the first fc layer.
+From Figure 21, we can see that removing the group penalty on the first fc layer leads to worse final performance.
+Furthermore, comparing the learned coefficients shows that they deviate more from the ground truth.
+Inspecting the weight matrix of the first fc layer, we also observe that the neural network begins to use the two redundant features, feature3 and feature4, even though they were not used in the data generation process for setting 1 (see D.2).
+D.10 Extra Results on PMLB Benchmark
+We conduct a more extensive evaluation of prediction performance using the PMLB benchmark Romano et al.
+(2021).
+We filter out datasets that were either too small (less than 1,000) or too large (more than 16,000), resulting in a total of 10 regression datasets.
+For each dataset, we compare the performance of Lasso, LassoNet, a standard Neural Network, and NIMO.
+Hyperparameters for all models are selected via grid search.
+The results are presented in Figure 22.
+Figure 22: Comparison of prediction performance on PMLB.
+As discussed in the limitations of NIMO (Sec.
+D.6), some datasets require additional feature engineering for NIMO to perform well.
+Within the PMLB benchmark, we indeed identified such a case.
+The 225_puma8NH dataset cannot be well fitted by NIMO directly (yielding an MSE of around 0.4).
+However, after applying a polynomial basis expansion, NIMO achieves performance comparable to LassoNet and Neural Networks.
 Appendix E LLM statement